### v2

# Correspondences between graph C*-algebras

Abstract.
We describe proper correspondences from graph $\textup{C}^{*}$-algebras to arbitrary $\textup{C}^{*}$-algebras by K-theoretic data. If the target $\textup{C}^{*}$-algebra is a graph $\textup{C}^{*}$-algebra as well, we may lift an isomorphism on a certain invariant to correspondences back and forth that are inverse to each other up to ideal-preserving homotopy. This implies a classification theorem for purely infinite graph $\textup{C}^{*}$-algebras up to stable isomorphism.
2010 Mathematics Subject Classification:
Primary 46L35; Secondary 19K35
Supported by the DFG grant “Classification of non-simple purely infinite C*-algebras.” This work is part of the project Graph Algebras partially supported by EU grant HORIZON-MSCA-SE-2021 Project 101086394.
1. Introduction
Graph $\textup{C}^{*}$-algebras form a class of $\textup{C}^{*}$-algebras that are well-behaved in various ways. For instance, they are all nuclear and satisfy the Universal Coefficient Theorem. Moreover, their ideal structure and K-theory is well-understood in terms of the underlying graph, and the K-theory has some special properties: the $\textup{K}_{1}$-group of a graph $\textup{C}^{*}$-algebra is always free, and the exponential map associated to a (gauge-invariant) ideal in a graph $\textup{C}^{*}$-algebra always vanishes. In addition, every simple graph $\textup{C}^{*}$-algebra is either approximately finite-dimensional or purely infinite. (See [Raeburn:Graph_algebras] for (references for) these well known facts.) Hence it is natural to explore what the general classification theory of $\textup{C}^{*}$-algebras has to say about $\textup{C}^{*}$-algebras with these features. Among other places, this has been done in [MR2562779, Eilers-Restorff-Ruiz:Strong_class_of_ext, Eilers-Tomforde:Graph, MR3056712, MR3194162, MR2666426] for the mixed cases and in [Meyer-Nest:Filtrated_K, Arklint-Bentmann-Katsura:Reduction, Bentmann:Real_rank_zero_and_int_cancellation, Bentmann-Meyer:More_general] for the (strongly) purely infinite case. While many results have been proved, a classification of nonunital graph $\textup{C}^{*}$-algebras along these lines is not yet in reach.
The approach to classification using the symbolic dynamics associated to the underlying graph has been very successful recently, culminating in the complete classification of unital graph $\textup{C}^{*}$-algebras (see [Eilers-Restorff-Ruiz-Sorensen:Complete] and the references therein, and [Restorff:Classification, Rordam:Class_of_CK] for earlier results). Unitality is, however, a strong restriction for graph $\textup{C}^{*}$-algebras because the unitisation of a graph $\textup{C}^{*}$-algebra need not be a graph $\textup{C}^{*}$-algebra. A graph $\textup{C}^{*}$-algebra is unital if and only if the underlying graph has finitely many vertices. It seems difficult to generalise the ideas of [Eilers-Restorff-Ruiz-Sorensen:Complete] to the nonunital case.
The present article introduces a new approach to classify graph $\textup{C}^{*}$-algebras. The goal is to understand (proper) $\textup{C}^{*}$-correspondences out of graph $\textup{C}^{*}$-algebras well enough to be able to establish existence and uniqueness results for proper $\textup{C}^{*}$-correspondences lifting isomorphisms on suitable invariants. We achieve this with uniqueness up to (ideal-preserving) homotopy. This is the main technical achievement of the article. Uniqueness up to homotopy is, however, too weak for Elliott’s Approximate Intertwining Argument. Hence our results do not suffice for a general classification of graph $\textup{C}^{*}$-algebras up to stable isomorphism. For purely infinite graph $\textup{C}^{*}$-algebras, however, it implies an equivalence in Kirchberg’s KK-theory, which is enough to apply Kirchberg’s Classification Theorem [Kirchberg:Michael]. So we get a classification result for all purely infinite (possibly nonunital) graph $\textup{C}^{*}$-algebras up to stable isomorphism. In general, our method only gives a classification theorem for graph $\textup{C}^{*}$-algebras up to ideal-preserving homotopy equivalence.
We now describe the sections of this article. In Section 2, we collect some basic results about proper $\textup{C}^{*}$-correspondences, their induced maps on ideal lattices and their relation to -homomorphisms into stabilised $\textup{C}^{*}$-algebras. Section 3 studies proper $\textup{C}^{*}$-correspondences out of graph $\textup{C}^{*}$-algebras. The main result says that any homomorphism between the monoids of equivalence classes of projections lifts to a proper $\textup{C}^{*}$-correspondence if the domain is a graph $\textup{C}^{*}$-algebra. Section 4 relates the monoid of equivalence classes of projections to the ideal lattice and the ordered $\textup{K}_{0}$-groups of ideals, especially in the case of graph $\textup{C}^{*}$-algebras. Section 5 discusses some homological algebra for certain diagrams of Abelian groups, especially the computation of the Ext groups, which is where the obstruction class and other homological invariants of correspondences live. Section 6 adds information related to $\textup{K}_{1}$ to the picture. We show that a pair of maps on equivalence classes of projections and on $\textup{K}_{1}$ is realised by a proper $\textup{C}^{*}$-correspondence if and only if a certain variant of the obstruction class of [Bentmann-Meyer:More_general] vanishes. In Section 7, we combine our new results with Kirchberg’s Classification Theorem to classify purely infinite graph $\textup{C}^{*}$-algebras. In Sections 8 and 9, we allow arbitrary graph $\textup{C}^{*}$-algebras again and prove an ideal-preserving homotopy equivalence instead of an isomorphism under the same assumptions as in Section 7. More precisely, Section 8 only proves a homotopy equivalence preserving gauge-invariant ideals. Section 9 contains some extra arguments that allow to prove an ideal-preserving homotopy equivalence for all ideals, gauge-invariant or not.
The authors thank the referee for carefully reading the article and spotting an error in an earlier version of this article.
2. Correspondences versus homomorphisms into stabilisations
This section contains results for general $\sigma$-unital $\textup{C}^{*}$-algebras, which we are going to specialise to graph $\textup{C}^{*}$-algebras in later sections. Let $A$ and $B$ be two $\textup{C}^{*}$-algebras. We are going to show that proper $\textup{C}^{*}$-correspondences between $A$ and $B$ are “equivalent” to -homomorphisms from $A$ to the $\textup{C}^{*}$-stabilisation $B\otimes\mathbb{K}$ when $A$ is $\sigma$-unital. Here $\mathbb{K}\mathrel{\vcentcolon=}\mathbb{K}(\ell^{2}\mathbb{N})$ is the $\textup{C}^{*}$-algebra of compact operators on a separable Hilbert space. If $A$ is $\textup{C}^{*}$-stable and both $A$ and $B$ are $\sigma$-unital, then full, proper correspondences are equivalent to nondegenerate -homomorphisms $A\to B\otimes\mathbb{K}$. Let $\mathbb{I}(A)$ be the lattice of closed ideals of $A$. A -homomorphism $\varphi\colon A\to B\otimes\mathbb{K}$ induces a map $\mathbb{I}(\varphi)\colon\mathbb{I}(A)\to\mathbb{I}(B)$. We describe it in terms of the proper $\textup{C}^{*}$-correspondence corresponding to $\varphi$. All this is either easy or follows from the proofs by Mingo and Phillips [Mingo-Phillips:Triviality] of the Kasparov Stabilisation Theorem and the Brown–Green–Rieffel Theorem.
First we recall some definitions. An $A,B$-correspondence is a (right) Hilbert $B$-module $\mathcal{H}$ with a nondegenerate -homomorphism $\varphi$ from $A$ to the $\textup{C}^{*}$-algebra $\mathbb{B}(\mathcal{H})$ of adjointable operators on $\mathcal{H}$. It is proper if $\varphi(A)$ is contained in the $\textup{C}^{*}$-algebra $\mathbb{K}(\mathcal{H})$ of compact operators on $\mathcal{H}$. Let $\langle\mathcal{H}\,|\,\mathopen{}\mathcal{H}\rangle\mathrel{\triangleleft}B$ be the ideal generated by the inner products $\langle x\,|\,\mathopen{}y\rangle$ for $x,y\in\mathcal{H}$. Call $\mathcal{H}$ full if $\langle\mathcal{H}\,|\,\mathopen{}\mathcal{H}\rangle=B$. An isomorphism between two $A,B$-correspondences $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ is a unitary operator $u\colon\mathcal{H}_{1}\xrightarrow{\sim}\mathcal{H}_{2}$ of right Hilbert $B$-modules that intertwines the left actions of $A$, that is, $u(a\cdot\xi)=a\cdot(u\xi)$ for all $a\in A$, $\xi\in\mathcal{H}_{1}$. The direct sum $\mathcal{H}_{1}\oplus\mathcal{H}_{2}$ of two correspondences is defined in the obvious way. It is again proper if $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are proper. And
(2.1) $\langle\mathcal{H}_{1}\oplus\mathcal{H}_{2}\,|\,\mathopen{}\mathcal{H}_{1}\oplus\mathcal{H}_{2}\rangle=\langle\mathcal{H}_{1}\,|\,\mathopen{}\mathcal{H}_{1}\rangle+\langle\mathcal{H}_{2}\,|\,\mathopen{}\mathcal{H}_{2}\rangle,$
so that $\mathcal{H}_{1}\oplus\mathcal{H}_{2}$ is full if $\mathcal{H}_{1}$ or $\mathcal{H}_{2}$ is full. The direct sum operation descends to isomorphism classes, is associative, and has the zero element $\{0\}$. Thus isomorphism classes of proper $A,B$-correspondences with the operation $\oplus$ induced by the direct sum of correspondences form a monoid. We denote it by $\mathcal{P}(A,B)$.
An $A,B$-correspondence $\mathcal{H}$ induces a map $\mathbb{I}(\mathcal{H})\colon\mathbb{I}(A)\to\mathbb{I}(B)$ by mapping $I\mathrel{\triangleleft}A$ to the ideal $\langle I\mathcal{H}\,|\,\mathopen{}I\mathcal{H}\rangle\mathrel{\triangleleft}B$ generated by the inner products $\langle ax\,|\,\mathopen{}by\rangle$ with $a,b\in I$, $x,y\in\mathcal{H}$. Since $I^{*}\cdot I=I$, we have
$\langle I\mathcal{H}\,|\,\mathopen{}I\mathcal{H}\rangle=\langle\mathcal{H}\,|\,\mathopen{}I\mathcal{H}\rangle=\langle I\mathcal{H}\,|\,\mathopen{}\mathcal{H}\rangle.$
Therefore, the map $\mathbb{I}(\mathcal{H})$ commutes with arbitrary suprema, that is,
$\mathbb{I}(\mathcal{H})\biggl{(}\bigvee_{I\in S}I\biggr{)}=\biggl{(}\bigvee_{I\in S}\mathbb{I}(\mathcal{H})(I)\biggr{)}$
for any set $S$ of ideals in $A$, where $\bigvee_{I\in S}I=\overline{\sum_{I\in S}I}$ is the supremum of $S$ in $\mathbb{I}(A)$. As a consequence, the map $\mathbb{I}(\mathcal{H})$ is monotone and preserves the zero ideals. It need not commute with intersections, however. And $\mathbb{I}(\mathcal{H})(A)=\langle\mathcal{H}\,|\,\mathopen{}\mathcal{H}\rangle$ is only equal to $B$ if $\mathcal{H}$ is full.
Let :$\mathcal{H}^{\infty}\mathrel{\vcentcolon=}\mathcal{H}\otimes\ell^{2}(\mathbb{N})$ or, equivalently, the direct sum of countably many copies of $\mathcal{H}$. There is an isomorphism $\mathbb{K}(\mathcal{H})\otimes\mathbb{K}\xrightarrow{\sim}\mathbb{K}(\mathcal{H}^{\infty})$, which maps $S\otimes T\in\mathbb{K}(\mathcal{H})\otimes\mathbb{K}$ to the unique operator $S\otimes T$ on $\mathcal{H}\otimes\ell^{2}\mathbb{N}$ that satisfies $(S\otimes T)(x\otimes y)\mathrel{\vcentcolon=}S(x)\otimes T(y)$ for all $x\in\mathcal{H}$, $y\in\ell^{2}\mathbb{N}$. In particular, $B\otimes\mathbb{K}\cong\mathbb{K}(B^{\infty})$ because $B\cong\mathbb{K}(B)$ when we view $B$ as a Hilbert $B$-module in the obvious way.
A -homomorphism $\varphi\colon A\to B\otimes\mathbb{K}$ induces a proper $A,B$-correspondence $\mathcal{H}_{\varphi}$ as follows. The underlying Hilbert $B$-module is :$\mathcal{H}\mathrel{\vcentcolon=}\varphi(A)\cdot B^{\infty}\subseteq B^{\infty}$, and $A$ acts nondegenerately on $\mathcal{H}$ through $\varphi$. If $\varphi$ is nondegenerate, then $\mathcal{H}=B^{\infty}$ and thus $\mathcal{H}_{\varphi}$ is full. The next proposition reverses this construction if $A$ is $\sigma$-unital, that is, if $A$ has a sequential approximate unit.
Proposition 2.2.
If $A$ is $\sigma$-unital, then any proper $A,B$-correspondence $\mathcal{H}$ is isomorphic to $\mathcal{H}_{\varphi}$ for a -homomorphism $\varphi\colon A\to B\otimes\mathbb{K}$. If $A$ and $B$ are $\sigma$-unital and $A$ is $\textup{C}^{*}$-stable, then any proper, full $A,B$-correspondence $\mathcal{H}$ is isomorphic to $\mathcal{H}_{\varphi}$ for a nondegenerate -homomorphism $\varphi\colon A\to B\otimes\mathbb{K}$.
Proof.
First assume only that $A$ is $\sigma$-unital. Let $\varphi_{0}\colon A\to\mathbb{K}(\mathcal{H})$ denote the left action on $\mathcal{H}$. Since it is nondegenerate, it maps a sequential approximate unit in $A$ to one in $\mathbb{K}(\mathcal{H})$. The Hilbert module $\mathcal{H}$ is countably generated if and only if $\mathbb{K}(\mathcal{H})$ is $\sigma$-unital. So $\mathcal{H}$ is countably generated. Then the Kasparov Stabilisation Theorem gives a unitary operator $\mathcal{H}\oplus B^{\infty}\cong B^{\infty}$ and, in particular, an adjointable isometry $S\colon\mathcal{H}\to B^{\infty}$. Then $\varphi(a)\mathrel{\vcentcolon=}S\varphi_{0}(a)S^{*}\in\mathbb{K}(B^{\infty})\cong B\otimes\mathbb{K}$ for $a\in A$ defines a -homomorphism $\varphi\colon A\to B\otimes\mathbb{K}$. By construction, $\varphi(A)\cdot B^{\infty}=S(\mathcal{H})$ and $S$ is an isomorphism of $A,B$-correspondences from $\mathcal{H}$ to $\mathcal{H}_{\varphi}$.
Now assume, in addition, that $A=A_{0}\otimes\mathbb{K}$ is $\textup{C}^{*}$-stable and that $B$ is $\sigma$-unital. The nondegenerate left action of $A$ on $\mathcal{H}$ induces an action of $\mathbb{K}\subseteq\mathcal{M}(A\otimes\mathbb{K})$. Let $p\in\mathbb{K}$ be a rank-$1$ projection and $\mathcal{H}_{0}\mathrel{\vcentcolon=}(1\otimes p)\mathcal{H}$. The matrix units in $\mathbb{K}$ allow to identify $\mathcal{H}\cong\mathcal{H}_{0}^{\infty}$. Since $\mathcal{H}$ is full, so is $\mathcal{H}_{0}$. Since $\mathcal{H}$ is countably generated, so is $\mathcal{H}_{0}$. Since $B=\mathbb{K}(B)$ is $\sigma$-unital, $B$ is a countably generated Hilbert $B$-module. Hence [Mingo-Phillips:Triviality]*Theorem 1.9 gives a unitary operator $S\colon\mathcal{H}\xrightarrow{\sim}\mathcal{H}_{0}^{\infty}\xrightarrow{\sim}B^{\infty}$. Now the -homomorphism $\varphi\colon A\to B\otimes\mathbb{K}$, $a\mapsto SaS^{*}$, is nondegenerate, and $S$ is an isomorphism from $\mathcal{H}$ to $\mathcal{H}_{\varphi}$. ∎
If $A=\mathbb{C}$, then a proper $\mathbb{C},B$-correspondence is just a Hilbert $B$-module $\mathcal{H}$ with $\textup{id}_{\mathcal{H}}\in\mathbb{K}(\mathcal{H})$. We call such Hilbert $B$-modules proper. Let $\mathbb{M}_{\infty}(B)=\bigcup_{n=0}^{\infty}\mathbb{M}_{n}(B)$ denote the set of infinite matrices over $B$ with finitely many nonzero entries.
Theorem 2.3.
Let $B$ be a $\textup{C}^{*}$-algebra. A Hilbert $B$-module $\mathcal{H}$ is proper if and only if there is a projection $p\in\mathbb{M}_{\infty}(B)$ with $\mathcal{H}\cong p\cdot B^{\infty}$. Let $p_{1},p_{2}\in\mathbb{M}_{\infty}(B)$ be projections. Unitary operators $p_{1}B^{\infty}\to p_{2}B^{\infty}$ are in bijection with Murray–von Neumann equivalences between $p_{1}$ and $p_{2}$, that is, $u\in\mathbb{M}_{\infty}(B)$ with $uu^{*}=p_{2}$ and $u^{*}u=p_{1}$.
Proof.
This follows from well known results if $B$ is unital (see [Wegge-Olsen:K-theory]*Theorem 15.4.2). The statements that we claim still hold if $B$ is nonunital (see also [Banerjee-Meyer:Coarse]*Proposition 4.2). We include a proof here for the convenience of the reader.
First assume that $\mathcal{H}\cong pB^{m}$ for a projection $p\in\mathbb{M}_{m}(B)$. We let $V\colon\mathcal{H}\to B^{m}$ be the resulting isometric embedding. It is adjointable with the adjoint $B^{m}\to\mathcal{H}$, $x\mapsto p\cdot x$. The compositions $V^{*}V$ and $VV^{*}$ are the identity map on $\mathcal{H}$ and left multiplication by $p$, respectively. The latter is a compact operator on $B^{m}$ because $\mathbb{K}(B^{m})=\mathbb{M}_{m}(B)$. Since $VV^{*}$ is compact, so is $V=VV^{*}V$ and hence also $V^{*}V=\textup{id}_{\mathcal{H}}$.
Conversely, assume that $\textup{id}_{\mathcal{H}}$ is compact. Then there is a finite-rank approximation $\sum_{j=1}^{m}{}\lvert x_{j}\rangle\langle y_{j}\rvert$ with, say,
$\bigg{\lVert}\sum_{j=1}^{m}{}\lvert x_{j}\rangle\langle y_{j}\rvert-\textup{id}_{\mathcal{H}}\bigg{\rVert}<1.$
Then $\sum_{j=1}^{m}{}\lvert x_{j}\rangle\langle y_{j}\rvert$ is invertible in $\mathbb{K}(\mathcal{H})$; let $s\in\mathbb{K}(\mathcal{H})$ be its inverse. Let
$y\mathrel{\vcentcolon=}(\langle y_{j}\rvert)_{j=1,\dotsc,m}\colon\mathcal{H}\to B^{m},\qquad x\mathrel{\vcentcolon=}(\lvert x_{j}\rangle)_{j=1,\dotsc,m}\colon B^{m}\to\mathcal{H};$
these are adjointable maps with $(s\circ x)\circ y=\textup{id}_{\mathcal{H}}$. Then $e=y\circ(s\circ x)\colon B^{m}\to B^{m}$ is idempotent, and adjointable as a composition of adjointable operators. Hence its range is complementable. The operators $e$ and $y$ have the same range and $y^{*}$ is surjective because $(s\circ x)\circ y=\textup{id}_{\mathcal{H}}$. As a consequence, the operator $y$ has a polar decomposition. This gives an isometry $V\colon\mathcal{H}\to B^{m}$. Since $\textup{id}_{\mathcal{H}}$ is compact, so is $V=V\circ\textup{id}_{\mathcal{H}}$ and hence $VV^{*}$. Thus $VV^{*}$ is left multiplication by some $p\in\mathbb{M}_{m}(B)$. This is a projection whose image is isomorphic to $\mathcal{H}$ via $V$. ∎
Thus $\mathcal{P}(B)\mathrel{\vcentcolon=}\mathcal{P}(\mathbb{C},B)$ is canonically isomorphic to the monoid of Murray–von Neumann equivalence classes of projections in $\mathbb{M}_{\infty}(B)$ with the usual operation $\oplus$, even if $B$ is not unital.
Proper correspondences may be composed by tensor product. This descends to a category structure on isomorphism classes of proper correspondences. In particular, there is a canonical composition $\mathcal{P}(A)\times\mathcal{P}(A,B)\to\mathcal{P}(B)$. It is additive in both variables. Equivalently, a proper $A,B$-correspondence $\mathcal{H}$ induces a monoid homomorphism $\mathcal{H}_{*}\colon\mathcal{P}(A)\to\mathcal{P}(B)$, $[\mathcal{M}]\mapsto[\mathcal{M}\otimes_{A}\mathcal{H}]$, and these homomorphisms satisfy $(\mathcal{H}_{1})_{*}+(\mathcal{H}_{2})_{*}=(\mathcal{H}_{1}\oplus\mathcal{H}_{2})_{*}$.
Next, we define “unitary equivalence” of -homomorphisms $A\to B\otimes\mathbb{K}$ in such a way that it corresponds exactly to unitary equivalence of the corresponding proper $A,B$-correspondences. Let $R_{\varphi}\mathrel{\vcentcolon=}\varphi(A)\cdot B\otimes\mathbb{K}$. This is a right ideal. We also view it as a Hilbert module over $B\otimes\mathbb{K}$.
Definition 2.4.
Let $\varphi_{0},\varphi_{1}\colon A\rightrightarrows B\otimes\mathbb{K}$ be two -homomorphisms. A unitary equivalence from $\varphi_{0}$ to $\varphi_{1}$ is a unitary operator $u\colon R_{\varphi_{0}}\xrightarrow{\sim}R_{\varphi_{1}}$ with $u\varphi_{0}(a)u^{*}=\varphi_{1}(a)$ for all $a\in A$, as operators on the Hilbert $B$-module $R_{\varphi_{1}}$.
A unitary multiplier $\bar{u}$ of $B\otimes\mathbb{K}$ with $u\varphi_{0}(a)u^{*}=\varphi_{1}(a)$ in $B\otimes\mathbb{K}$ for all $a\in A$ restricts to a unitary operator $u\colon R_{\varphi_{0}}\xrightarrow{\sim}R_{\varphi_{1}}$ that is a unitary equivalence from $\varphi_{0}$ to $\varphi_{1}$. Thus we allow more unitary equivalences than usual.
The argument that identifies the multiplier algebra of $\mathbb{K}(\mathcal{H})$ for a Hilbert $B$-module $\mathcal{H}$ with $\mathbb{B}(\mathcal{H})$ gives a bijection between unitary equivalences $\varphi_{1}\to\varphi_{2}$ and isomorphisms of correspondences $\mathcal{H}_{\varphi_{1}}\to\mathcal{H}_{\varphi_{2}}$. Thus isomorphism classes of proper $A,B$-correspondences are in bijection with unitary equivalence classes of -homomorphisms $A\to B\otimes\mathbb{K}$. And isomorphism classes of proper, full $A\otimes\mathbb{K},B$-correspondences are in bijection with unitary equivalence classes of nondegenerate -homomorphisms $A\otimes\mathbb{K}\to B\otimes\mathbb{K}$, for the standard notion of unitary equivalence.
A -homomorphism $\varphi\colon A\to B\otimes\mathbb{K}$ induces a map $\mathbb{I}(\varphi)\colon\mathbb{I}(A)\to\mathbb{I}(B\otimes\mathbb{K})\cong\mathbb{I}(B)$ by mapping $I\mathrel{\triangleleft}A$ to the ideal in $B\otimes\mathbb{K}$ generated by $\varphi(I)\subseteq B\otimes\mathbb{K}$. A routine computation identifies this map with the map $\mathbb{I}(\mathcal{H}_{\varphi})$ defined above.
For $\textup{C}^{*}$-algebra classification, the equivalence relation of unitary equivalence for -homomorphisms is too fine, even with our mild modification. We should use asymptotic or approximate unitary equivalence or homotopy. A homotopy from $\varphi_{0}$ to $\varphi_{1}$ is a -homomorphism $\Phi\colon A\to\mathrm{C}([0,1],B\otimes\mathbb{K})$ with $\textup{ev}_{t}\circ\Phi=\varphi_{t}$ for $t=0,1$. An asymptotic unitary equivalence from $\varphi_{0}$ to $\varphi_{1}$ is a family of unitary operators $u_{t}\colon R_{\varphi_{0}}\xrightarrow{\sim}R_{\varphi_{1}}$ for $t\in[0,\infty)$ that is -strongly continuous in the sense that $t\mapsto u_{t}(x)\in R_{\varphi_{1}}$ and $t\mapsto u_{t}^{*}(y)\in R_{\varphi_{0}}$ are norm-continuous for $x\in R_{\varphi_{0}}$, $y\in R_{\varphi_{1}}$ and that satisfies
$\lim_{t\to\infty}u_{t}\varphi_{0}(a)u_{t}^{*}=\varphi_{1}(a)$
for all $a\in A$, as operators on the Hilbert $B$-module $R_{\varphi_{1}}$. An approximate unitary equivalence from $\varphi_{0}$ to $\varphi_{1}$ is a family of unitary operators $u_{n}\colon R_{\varphi_{0}}\xrightarrow{\sim}R_{\varphi_{1}}$ for $n\in\mathbb{N}$ that satisfies
$\lim_{n\to\infty}u_{n}\varphi_{0}(a)u_{n}^{*}=\varphi_{1}(a)$
for all $a\in A$. For proper $A,B$-correspondences, these three notions correspond to the following:
Definition 2.5.
A homotopy from $\mathcal{H}_{0}$ to $\mathcal{H}_{1}$ is an $A,\mathrm{C}([0,1],B)$-correspondence $\mathcal{H}$ together with isomorphisms $\mathcal{H}\otimes_{\textup{ev}_{t}}B\cong\mathcal{H}_{t}$ for $t=0,1$. An asymptotic isomorphism from $\mathcal{H}_{0}$ to $\mathcal{H}_{1}$ is a -strongly continuous family of unitary operators $(u_{t})_{t\in[0,\infty)}$ between the underlying Hilbert $B$-modules that satisfies
$\lim_{t\to\infty}u_{t}(a\cdot_{0}u_{t}^{*}x)=a\cdot_{1}x$
for all $a\in A$, $x\in\mathcal{H}_{1}$; here $\cdot_{i}$ denotes the left action of $A$ on $\mathcal{H}_{i}$. An approximate isomorphism is defined similarly, replacing the indexing set $[0,\infty)$ by $\mathbb{N}$.
3. Correspondences out of a graph $C^{*}$-algebra
Let $G=(r,s\colon E\rightrightarrows V)$ be a directed graph with sets of vertices and edges $V$ and $E$, respectively, and range and source maps $r$ and $s$. The graph $\textup{C}^{*}$-algebra of $G$ is the universal $\textup{C}^{*}$-algebra $\textup{C}^{*}(G)$ generated by mutually orthogonal projections $p_{v}$ for $v\in V$ and partial isometries $t_{e}$ for $e\in E$ subject to the relations
  1. (CK1)
    $t_{e}^{*}t_{e}=p_{s(e)}$ for all $e\in E$;
  2. (CK2)
    $\sum_{r(e)=v}t_{e}t_{e}^{*}=p_{v}$ for all $v\in V$ with $0<\lvert r^{-1}(v)\rvert<\infty$;
  3. (CK3)
    $t_{e}t_{e}^{*}\leq p_{r(e)}$ for all $e\in E$ and $t_{e}^{*}t_{f}=0$ for all $e,f\in E$ with $e\neq f$.
We call $v\in V$ regular if $0<\lvert r^{-1}(v)\rvert<\infty$; let $V_{\mathrm{reg}}\subseteq V$ be the subset of regular vertices. The graph $G$ is regular if all vertices are regular. Then the relation (CK2) is imposed for all vertices. More generally, the graph $G$ is row-finite if $\lvert r^{-1}(v)\rvert<\infty$ for all $v\in V$. Then the relation (CK3) is redundant. Any graph $\textup{C}^{*}$-algebra $\textup{C}^{*}(G)$ is stably isomorphic to $\textup{C}^{*}(G^{\prime})$ for a regular graph $G^{\prime}$, which is constructed in [Muhly-Tomforde:Adding_tails] by “adding tails” to $G$. Since we only care about graph $\textup{C}^{*}$-algebras up to stable isomorphism, we may as well assume $G$ to be regular. Actually, our methods work equally well for row-finite graphs. We assume all graphs in the following to be row-finite, sometimes tacitly.
Theorem 3.1 ([Ara-Moreno-Pardo:Nonstable_K_for_Graph_Algs]*Theorems 3.5 and 7.1).
Let $G$ be a row-finite graph. Each projection $q\in\mathbb{M}_{\infty}(\textup{C}^{*}(G))$ is equivalent to a finite sum $\sum_{v\in V}c_{v}p_{v}$ of vertex projections with $(c_{v})_{v\in V}\in\mathbb{N}[V]$, that is, $c_{v}\in\mathbb{N}$ for all $v\in V$ and $c_{v}\neq 0$ only for finitely many $v\in V$.
Theorem 3.2.
Let $G$ be a row-finite graph and $B$ a $\textup{C}^{*}$-algebra. Proper $\textup{C}^{*}(G),B$-correspondences are equivalent to families $\bigl{(}(\mathcal{H}_{v})_{v\in V},(U_{v})_{v\in V_{\mathrm{reg}}}\bigr{)}$, where each $\mathcal{H}_{v}$ is a proper Hilbert $B$-module and $U_{v}$ is a unitary operator $\bigoplus_{e\in E^{v}}\mathcal{H}_{s(e)}\xrightarrow{\sim}\mathcal{H}_{v}$. Here “equivalent” means an equivalence of categories, where arrows from $\bigl{(}(\mathcal{H}_{1,v}),(U_{1,v})\bigr{)}$ to $\bigl{(}(\mathcal{H}_{2,v}),(U_{2,v})\bigr{)}$ are families of unitaries $W_{v}\colon\mathcal{H}_{1,v}\xrightarrow{\sim}\mathcal{H}_{2,v}$ for $v\in V$ that make the following diagrams commute for all $v\in V_{\mathrm{reg}}$:
[…]
For any monoid homomorphism $\varphi\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$, there is a proper $\textup{C}^{*}(G),B$-correspondence $\mathcal{H}$ with $\mathcal{H}_{*}=\varphi$.
Proof.
Fix $v\in V_{\mathrm{reg}}$. Let $E^{v}\mathrel{\vcentcolon=}\{e\in E\mid r(e)=v\}$ and let $m_{v}=\lvert E^{v}\rvert$ be the number of edges with range $v$. Totally order these edges, and let $u_{v}\in\mathbb{M}_{1,m_{v}}(\textup{C}^{*}(G))$ be the row matrix with entries $t_{e}$ for $e\in E^{v}$. The relations (CK1) and (CK2) say that $u_{v}u_{v}^{*}=\sum_{e\in E^{v}}t_{e}t_{e}^{*}=p_{v}$ and $u_{v}^{*}u_{v}=(t_{e}^{*}t_{f})_{e,f\in E^{v}}=(\delta_{e,f}p_{s(e)})_{e,f\in E^{v}}$, the orthogonal direct sum $\bigoplus_{e\in E^{v}}p_{s(e)}$ of the projections $p_{s(e)}$ for $e\in E^{v}$. That is, $u_{v}$ is a partial isometry with range and source projections $p_{v}$ and $\bigoplus_{e\in E^{v}}p_{s(e)}$. Since any vertex in $r(E)$ is regular, the unitaries $u_{v}$ for $v\in V_{\mathrm{reg}}$ contain the partial isometries $t_{e}$ for all edges.
Now we prove that a proper $\textup{C}^{*}(G),B$-correspondence $\mathcal{H}$ is equivalent to a family $\bigl{(}(\mathcal{H}_{v})_{v\in V},(U_{v})_{v\in V_{\mathrm{reg}}}\bigr{)}$. Let $\mathcal{H}_{v}\mathrel{\vcentcolon=}p_{v}\cdot\mathcal{H}\subseteq\mathcal{H}$. These are proper Hilbert $B$-modules because $p_{v}\in\mathbb{K}(\mathcal{H})$ for a proper $\textup{C}^{*}(G),B$-correspondence. The partial isometry $u_{v}$ for $v\in V_{\mathrm{reg}}$ acts by a unitary operator $U_{v}\colon\bigoplus_{e\in E^{v}}\mathcal{H}_{s(e)}\xrightarrow{\sim}\mathcal{H}_{v}$. This maps $\mathcal{H}$ to a family $((\mathcal{H}_{v}),(U_{v}))$. Let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be proper $\textup{C}^{*}(G),B$-correspondences and let $W\colon\mathcal{H}_{1}\to\mathcal{H}_{2}$ be an isomorphism. This restricts to isomorphisms of Hilbert $B$-modules $W_{v}\colon\mathcal{H}_{1,v}\to\mathcal{H}_{2,v}$ because $W$ intertwines the left action of the projections $p_{v}$. The diagrams (3.3) commute because $W$ intertwines the left action of the generators $t_{e}$ for $e\in E$ and hence of the matrices $u_{v}$ for $v\in V_{\mathrm{reg}}$.
Conversely, let $\mathcal{H}_{v}$ for $v\in V$ be proper Hilbert $B$-modules and let
$U_{v}\colon\bigoplus_{e\in E^{v}}\mathcal{H}_{s(e)}\xrightarrow{\sim}\mathcal{H}_{v}$
be unitaries for $v\in V_{\mathrm{reg}}$. Let :$\mathcal{H}\mathrel{\vcentcolon=}\bigoplus_{v\in V}\mathcal{H}_{v}$. Let $P_{v}\in\mathbb{K}(\mathcal{H})$ be the projection onto the summand $\mathcal{H}_{v}$. Let $T_{e}$ for $e\in E^{v}$ be the restriction of $U_{v}$ to the summand $\mathcal{H}_{s(e)}$. The operators $P_{v}$ and $T_{e}$ on $\mathcal{H}$ satisfy the Cuntz–Krieger relations (CK1) and (CK2). So there is a unique -homomorphism $\textup{C}^{*}(G)\to\mathbb{K}(\mathcal{H})$ mapping $p_{v}$ to $P_{v}$ and $t_{e}$ to $T_{e}$. This is nondegenerate because it maps $\bigl{(}\bigoplus_{v\in F}p_{v}\bigr{)}_{F}$ to an approximate unit in $\mathbb{K}(\mathcal{H})$. Thus the family $((\mathcal{H}_{v}),(U_{v}))$ comes from a proper $\textup{C}^{*}(G),B$-correspondence. Let $((\mathcal{H}_{1,v}),(U_{1,v}))$ and $((\mathcal{H}_{2,v}),(U_{2,v}))$ be two such families and let $(W_{v})_{v\in V}$ be a family of unitaries $W_{v}\colon\mathcal{H}_{1,v}\xrightarrow{\sim}\mathcal{H}_{2,v}$ making the diagrams (3.3) commute for all $v\in V_{\mathrm{reg}}$. Then $W\mathrel{\vcentcolon=}\bigoplus_{v\in V}W_{v}$ is a unitary $\bigoplus_{v\in V}\mathcal{H}_{1,v}\to\bigoplus_{v\in V}\mathcal{H}_{2,v}$, which intertwines the left actions of the generators $p_{v}$ for $v\in V$ and $t_{e}$ for $e\in E$.
The functors from the category of proper $\textup{C}^{*}(G),B$-correspondences to the category of families $((\mathcal{H}_{v}),(U_{v}))$ and back are easily seen to be inverse to each other up to natural isomorphism. That is, they form an equivalence of categories.
Now let $\varphi\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$ be a homomorphism. We construct a family $((\mathcal{H}_{v}),(U_{v}))$ from it. For $v\in V$, let $\mathcal{H}_{v}$ be a proper Hilbert $B$-module that represents $\varphi[p_{v}]\in\mathcal{P}(B)$. Let $v\in V_{\mathrm{reg}}$. The equivalence $u_{v}$ in $\mathbb{M}_{\infty}(\textup{C}^{*}(G))$ implies that $\sum_{e\in E^{v}}{}[p_{s(e)}]=[p_{v}]$ holds in $\mathcal{P}(\textup{C}^{*}(G))$. Since $\varphi$ is a homomorphism, this implies $\sum_{e\in E^{v}}\varphi[p_{s(e)}]=\varphi[p_{v}]$. Equivalently, there is a unitary $U_{v}$ between the corresponding proper Hilbert $B$-modules $\bigoplus_{e\in E^{v}}\mathcal{H}_{s(e)}$ and $\mathcal{H}_{v}$. The Hilbert modules $\mathcal{H}_{v}$ for $v\in V$ and the unitaries $U_{v}$ for $v\in V_{\mathrm{reg}}$ describe a proper $\textup{C}^{*}(G),B$-correspondence $\mathcal{H}$, which induces a map $\mathcal{H}_{*}\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$. We claim that $\mathcal{H}_{*}=\varphi$. By construction, $\mathcal{H}_{*}[p_{v}]=[\mathcal{H}_{v}]=\varphi[p_{v}]$ for all $v\in V$. The subset of $\mathcal{P}(\textup{C}^{*}(G))$ where $\mathcal{H}_{*}$ and $\varphi$ coincide is a submonoid. Hence it contains all finite sums of the vertex projections. Any class in $\mathcal{P}(\textup{C}^{*}(G))$ is such a finite sum by Theorem 3.1. Hence $\mathcal{H}_{*}=\varphi$ holds on all of $\mathcal{P}(\textup{C}^{*}(G))$. ∎
Remark 3.4.
By Theorem 2.3, any proper Hilbert $B$-module is isomorphic to $p\cdot B^{\infty}$ for some projection $p\in\mathbb{M}_{\infty}(B)$, and unitary operators between proper Hilbert $B$-modules are equivalent to equivalences of projections. Hence we may replace the families $\bigl{(}(\mathcal{H}_{v})_{v\in V},(U_{v})_{v\in V_{\mathrm{reg}}}\bigr{)}$ by families $\bigl{(}(P_{v})_{v\in V},(U_{v})_{v\in V_{\mathrm{reg}}}\bigr{)}$ consisting of projections in $\mathbb{M}_{\infty}(B)$ and equivalences between them.
Let $\varphi\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$ be a homomorphism. There may be several proper $\textup{C}^{*}(G),B$-correspondences $\mathcal{H}$ with $\mathcal{H}_{*}=\varphi$. By Theorem 3.2, $\mathcal{H}$ is given by proper Hilbert $B$-modules $\mathcal{H}_{v}$ for $v\in V$ and unitaries $U_{v}\colon\bigoplus_{e\in E^{v}}\mathcal{H}_{s(e)}\to\mathcal{H}_{v}$ for $v\in V_{\mathrm{reg}}$. We must have $[\mathcal{H}_{v}]=\varphi[p_{v}]$ in $\mathcal{P}(B)$, which determines $\mathcal{H}_{v}$ up to isomorphism. Given any family of Hilbert $B$-module isomorphisms $W_{v}\colon\mathcal{H}_{v}\to\mathcal{H}_{v}^{\prime}$ for $v\in V$, there are unique unitaries $U_{v}^{\prime}$ for $v\in V_{\mathrm{reg}}$ that make (3.3) commute, so that $(W_{v})$ induces an isomorphism $(\mathcal{H}_{v},U_{v})\cong(\mathcal{H}^{\prime}_{v},U^{\prime}_{v})$. Thus it does not matter how we choose the Hilbert $B$-modules $\mathcal{H}_{v}$ in their isomorphism classes. But the choice of the operators $U_{v}$ does matter. They contain information about the action of the $\textup{C}^{*}$-correspondence on $\textup{K}_{1}$. We shall make this more precise later. Here we prepare this with some simple observations.
Lemma 3.5.
Let $\mathcal{H}_{v}$ for $v\in V$ be Hilbert $B$-modules. Let $U_{v}\colon\bigoplus_{e\in E^{v}}\mathcal{H}_{s(e)}\xrightarrow{\sim}\mathcal{H}_{v}$ for $v\in V_{\mathrm{reg}}$ be unitaries. Families of unitaries $U_{v}^{\prime}\colon\bigoplus_{e\in E^{v}}\mathcal{H}_{s(e)}\xrightarrow{\sim}\mathcal{H}_{v}$ for $v\in V_{\mathrm{reg}}$ are in bijection with families of unitaries $\Upsilon_{v}\in\mathcal{U}(\mathcal{H}_{v})$ for $v\in V_{\mathrm{reg}}$, by mapping $U^{\prime}_{v}$ to $\Upsilon_{v}\mathrel{\vcentcolon=}U^{\prime}_{v}U_{v}^{*}\in\mathcal{U}(\mathcal{H}_{v})$ and $\Upsilon_{v}\in\mathcal{U}(\mathcal{H}_{v})$ to $U^{\prime}_{v}\mathrel{\vcentcolon=}\Upsilon_{v}U_{v}$. An isomorphism between the $\textup{C}^{*}(G),B$-correspondences described by $((\mathcal{H}_{v}),(U_{v}))$ and $((\mathcal{H}_{v}),(\Upsilon_{v}U_{v}))$ is a family of unitaries $W_{v}\in\mathcal{U}(\mathcal{H}_{v})$ for $v\in V$ with
$\Upsilon_{v}=W_{v}\circ U_{v}\circ\biggl{(}\bigoplus_{e\in E^{v}}W_{s(e)}^{*}\biggr{)}\circ U_{v}^{*}$
for all $v\in V_{\mathrm{reg}}$.
Proof.
The bijective correspondence between $U^{\prime}_{v}$ and $\Upsilon_{v}$ is trivial. The condition $\Upsilon_{v}=W_{v}\circ U_{v}\circ\left(\bigoplus_{e\in E^{v}}W_{s(e)}\right)^{*}\circ U_{v}^{*}$ is equivalent to the commutativity of the diagram (3.3) with $U_{1,v}=U_{v}$ and $U_{2,v}=\Upsilon_{v}U_{v}$. ∎
Lemma 3.5 allows us to describe homotopies and asymptotic and approximate isomorphisms between $\textup{C}^{*}(G),B$-correspondences. Since homotopic projections are equivalent, two correspondences that are homotopic must have isomorphic Hilbert modules $\mathcal{H}_{v}$. The same holds for asymptotically or approximately isomorphic correspondences. Assume for simplicity that they even have the same Hilbert modules $\mathcal{H}_{v}$. By Lemma 3.5, the two correspondences are described by unitaries $U_{v},\Upsilon_{v}U_{v}\colon\bigoplus_{e\in E^{v}}\mathcal{H}_{s(e)}\rightrightarrows\mathcal{H}_{v}$ for some $\Upsilon_{v}\in\mathcal{U}(\mathcal{H}_{v})$ for $v\in V_{\mathrm{reg}}$, respectively. A homotopy between them consists of norm-continuous homotopies $(\Upsilon_{v,t})_{t\in[0,1]}$ in $\mathcal{U}(\mathcal{H}_{v})$ for all $v\in V_{\mathrm{reg}}$ and unitaries $W_{v,j}\in\mathcal{U}(\mathcal{H}_{v})$ for $j=0,1$ and $v\in V$, such that
$\Upsilon_{v,0}$ $=W_{v,0}\circ U_{v}\circ\biggl{(}\bigoplus_{e\in E^{v}}W_{s(e),0}^{*}\biggr{)}\circ U_{v}^{*},$
$\Upsilon_{v,1}$ $=W_{v,1}\circ\Upsilon_{v}U_{v}\circ\biggl{(}\bigoplus_{e\in E^{v}}W_{s(e),1}^{*}\biggr{)}\circ U_{v}^{*}.$
In other words, $(\Upsilon_{v,t})_{v\in V_{\mathrm{reg}}}$ for $t\in[0,1]$ yields a homotopy between two correspondences that are isomorphic to those corresponding to $(U_{v})_{v\in V_{\mathrm{reg}}}$ and $(\Upsilon_{v}U_{v})_{v\in V_{\mathrm{reg}}}$.
An asymptotic isomorphism between the two correspondences is given by a family of norm-continuous paths of $W_{v,t}\in\mathcal{U}(\mathcal{H}_{v})$ for $t\in[0,\infty\mathclose{[}$, $v\in V$, such that
$\lim_{t\to\infty}W_{v,t}\circ U_{v}\circ\biggl{(}\bigoplus_{e\in E^{v}}W_{s(e),t}^{*}\biggr{)}\circ U_{v}^{*}=\Upsilon_{v}$
for all $v\in V_{\mathrm{reg}}$; there is no constraint on $W_{v,0}$. An approximate isomorphism between them is similar, but instead $W_{v,t}\in\mathcal{U}(\mathcal{H}_{v})$ is needed only for $t\in\mathbb{N}$, $v\in V$. We can make this even more explicit:
Lemma 3.6.
Let $V$ be countable and let $((\mathcal{H}_{v}),(U_{v}))$ and $((\mathcal{H}_{v}),(\Upsilon_{v}U_{v}))$ describe two $\textup{C}^{*}(G),B$-correspondences. They are approximately isomorphic if and only if for each finite subset $F\subseteq V$ and each $\varepsilon>0$ there is a family of unitaries $W_{v}\in\mathcal{U}(\mathcal{H}_{v})$ for $v\in V$ with
$\left\lVert\Upsilon_{v}-W_{v}\circ U_{v}\circ\biggl{(}\bigoplus_{e\in E^{v}}W_{s(e)}^{*}\biggr{)}\circ U_{v}^{*}\right\rVert<\varepsilon$
for all $v\in F\cap V_{\mathrm{reg}}$.
Proof.
Write $V=\bigcup_{n\in\mathbb{N}}F_{n}$ as an increasing union of finite subsets. Let the unitaries $W_{v,n}$ be as in the statement of the lemma for $F_{n}\subseteq V$ and $\varepsilon=2^{-n}$. These unitaries give the desired approximate isomorphism. ∎
3.1. The AF-case
In this short subsection, we show that the techniques above already suffice to classify $\textup{C}^{*}$-correspondences from $\textup{C}^{*}(G)$ to $B$ if our graph $\textup{C}^{*}$-algebra is AF. This should be expected because the following sections are mainly about incorporating $\textup{K}_{1}(\textup{C}^{*}(G))$ into the picture, which is trivial in the AF-case.
Definition 3.7.
A cycle in a directed graph $G$ is a closed path $e_{1},\dotsc,e_{n}$, that is, $s(e_{1})=r(e_{n})$, $s(e_{2})=r(e_{1})$, …, $s(e_{n})=r(e_{n-1})$. A cycle is simple if the vertices $s(e_{1}),\dotsc,s(e_{n})$ are all distinct.
Assume for some time that $G$ has no cycles. Equivalently, $\textup{C}^{*}(G)$ is an AF-algebra. Then the invariant $\mathcal{P}(\textup{C}^{*}(G))$ determines $\textup{C}^{*}(G)$ up to isomorphism. We are going to prove that proper $\textup{C}^{*}(G),B$-correspondences are classified by the induced map $\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$ up to approximate isomorphism:
Proposition 3.8.
Let $G$ be a directed graph without cycles. Let $B$ be a $\textup{C}^{*}$-algebra. Identify proper $\textup{C}^{*}(G),B$-correspondences with families $((\mathcal{H}_{v})_{v\in V},(U_{v})_{v\in V_{\mathrm{reg}}})$. Assume that $((\mathcal{H}_{v}),(U_{v}))$ and $((\mathcal{H}^{\prime}_{v}),(U^{\prime}_{v}))$ induce the same map $\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$.
  1. (1)
    Let $F\subseteq V_{\mathrm{reg}}$ be a finite subset. Then $((\mathcal{H}^{\prime}_{v}),(U^{\prime}_{v}))$ is isomorphic to a proper $\textup{C}^{*}(G),B$-correspondence $((\mathcal{H}^{\prime\prime}_{v}),(U^{\prime\prime}_{v}))$ with $\mathcal{H}^{\prime\prime}_{v}=\mathcal{H}_{v}$ for all $v\in V$ and $U^{\prime\prime}_{v}=U_{v}$ for $v\in F$.
  2. (2)
    If $G$ is countable, then $((\mathcal{H}_{v}),(U_{v}))$ and $((\mathcal{H}^{\prime}_{v}),(U^{\prime}_{v}))$ are approximately isomorphic.
Proof.
We have already explained how to arrange $\mathcal{H}^{\prime}_{v}=\mathcal{H}_{v}$ for all $v\in V$. Define $x\prec y$ for $x,y\in V$ if there is a path from $x$ to $y$ in $G$. This is a partial order on $V$ because $G$ has no cycles. Let $F\subseteq V_{\mathrm{reg}}$ be finite. We may enumerate $F\cong\{1,\dotsc,n\}$ so that $x\prec y$ implies $x\leq y$ in $\{1,\dotsc,n\}$. We construct the isomorphism from $((\mathcal{H}^{\prime}_{v}),(U^{\prime}_{v}))$ to $((\mathcal{H}^{\prime\prime}_{v}),(U^{\prime\prime}_{v}))$ by induction on $n$, the number of points in $F$. We assume that we have such an isomorphism for $\{1,\dotsc,n-1\}$. Since isomorphisms of correspondences may be composed, we may assume that already $U_{i}^{\prime}=U_{i}$ for $i=1,\dotsc,n-1$. Now we define an isomorphism $(W_{v})_{v\in V}$ by $W_{n}\mathrel{\vcentcolon=}U_{n}(U_{n}^{\prime})^{*}$ and $W_{v}\mathrel{\vcentcolon=}1$ for all $v\neq n$ and, in particular, for all $v\in V\setminus F$. By Lemma 3.6, this is an isomorphism from $((\mathcal{H}^{\prime}_{v}),(U^{\prime}_{v}))$ to $((\mathcal{H}^{\prime}_{v}),(U^{\prime\prime}_{v}))$ with
$U_{v}^{\prime\prime}\mathrel{\vcentcolon=}W_{v}\circ U_{v}^{\prime}\circ\biggl{(}\bigoplus_{r(e)=v}W_{s(e)}^{*}\biggr{)}.$
If $v<n$, then $s(e)<n$ for all $e\in E$ with $r(e)=v$, so that $U_{v}^{\prime\prime}=U_{v}^{\prime}=U_{v}$. If $v=n$, then $s(e)<n$ for all $e\in E$ with $r(e)=n$, so that $U_{n}^{\prime\prime}=W_{n}U_{n}^{\prime}=U_{n}$. This proves the first statement. If $G$ is countable, we may enumerate $V_{\mathrm{reg}}\cong\mathbb{N}$. The isomorphisms $((\mathcal{H}_{v}),(U_{v}^{\prime}))\cong((\mathcal{H}_{v}),(U_{v}^{\prime\prime}))$ for the finite subsets $\{1,\dotsc,n\}\subseteq V$ form an approximate isomorphism between the $\textup{C}^{*}(G),B$-correspondences $((\mathcal{H}_{v}),(U_{v}^{\prime}))$ and $((\mathcal{H}_{v}),(U_{v}))$ because -polynomials in the generators $p_{v},s_{e}$ are dense in $\textup{C}^{*}(G)$. So the first statement implies the second one. ∎
4. Projections versus K-theory diagrams
For a $\textup{C}^{*}$-algebra $B$ that satisfies “stable weak cancellation,” we are going to describe $\mathcal{P}(B)$ using the ideal structure of $B$ and the $\textup{K}_{0}^{+}$-monoids for certain ideals of $B$. Graph $\textup{C}^{*}$-algebras and purely infinite $\textup{C}^{*}$-algebras have stable weak cancellation. This clarifies the relationship between $\mathcal{P}(B)$ and invariants that have already been used in the classification of nonsimple $\textup{C}^{*}$-algebras.
Let $p\in\mathbb{M}_{\infty}(B)$ be a projection. The ideal $\operatorname{supp}p$ generated by $p$ is the closed ideal generated by products $b_{1}^{*}pb_{2}$ with $b_{1},b_{2}\in B^{\infty}$. This is equal to the ideal generated by inner products of vectors in the Hilbert module $p\cdot B^{\infty}$ and to the ideal generated by the matrix coefficients of $p$. Any projection $p\in\mathcal{P}(B)$ belongs to $\mathbb{M}_{\infty}(\operatorname{supp}p)$ and is a full projection there by construction of $\operatorname{supp}p$. Equivalent projections generate the same ideal. Hence we get a map
$\operatorname{supp}\colon\mathcal{P}(B)\to\mathbb{I}(B).$
It is additive, that is, $\operatorname{supp}0=0$ and
(4.1) $\operatorname{supp}(p\oplus q)=\operatorname{supp}p+\operatorname{supp}q$
for all $p,q\in\mathcal{P}(B)$. Hence it is also monotone:
(4.2) $p,q\in\mathcal{P}(B),\ p\leq q\quad\Longrightarrow\quad\operatorname{supp}p\subseteq\operatorname{supp}q;$
here $p\leq q$ by definition if there is $p^{\prime}\in\mathcal{P}(B)$ with $p\oplus p^{\prime}=q$.
Definition 4.3.
Let $\mathbb{I}_{\mathrm{p}}(B)\subseteq\mathbb{I}(B)$ be the range of the map $\operatorname{supp}$. We also say that an ideal is generated by a projection if it is equal to $\operatorname{supp}p$ for some projection $p\in\mathbb{M}_{\infty}(B)$. Let $\mathbb{I}_{\mathrm{p}}^{\infty}(B)\subseteq\mathbb{I}(B)$ be the larger subset of all ideals in $B$ that are generated by some set of projections in $\mathbb{M}_{\infty}(B)$.
The lattice $\mathbb{I}(B)$ is isomorphic to the lattice of open subsets of the topological space $\widehat{B}$ of irreducible representations of $B$ (see [Dixmier:Cstar-algebres]*§3.2). Let $\widehat{I}\subseteq\widehat{B}$ denote the open subset of $\widehat{B}$ corresponding to $I\in\mathbb{I}(B)$. This space is indeed homeomorphic to the space of irreducible representations of $I$.
Lemma 4.4.
Let $B$ be a $\textup{C}^{*}$-algebra and $I\in\mathbb{I}(B)$. Then $I\in\mathbb{I}_{\mathrm{p}}(B)$ if and only if $I\in\mathbb{I}_{\mathrm{p}}^{\infty}(B)$ and $\widehat{I}$ is quasi-compact.
Proof.
Let $p\in\mathbb{M}_{n}(B)$ for some $n\in\mathbb{N}$ be idempotent. Let $I$ be the ideal generated by $p$. An irreducible representation $\varrho$ of $B$ remains nondegenerate on $I$ if and only if $\varrho_{n}(p)\neq 0$, if and only if $\lVert\varrho_{n}(p)\rVert\geq 1$, where $\varrho_{n}$ is the corresponding irreducible representation of $\mathbb{M}_{n}(B)$. So $\widehat{I}\subseteq\widehat{B}$ is the set of all $[\varrho]$ with $\lVert\varrho_{n}(p)\rVert\geq 1$. This subset is quasi-compact by [Dixmier:Cstar-algebres]*Proposition 3.3.7. Conversely, let $I$ be generated by some set of projections $(p_{\alpha})_{\alpha\in S}$ and let $\widehat{I}\subseteq\widehat{B}$ be quasi-compact. For each finite subset $F\subseteq S$, let $I_{F}$ be the ideal generated by the projections $p_{\alpha}$ for $\alpha\in F$. Then $\bigcup I_{F}=I$. Since $\widehat{I}$ is quasi-compact, there is some finite subset $F\subseteq S$ with $\widehat{I_{F}}=\widehat{I}$ and hence $I_{F}=I$. Thus $I$ is generated by the projection $\bigoplus_{\alpha\in F}p_{\alpha}$. ∎
Remark 4.5.
A $\textup{C}^{*}$-algebra has the ideal property if $\mathbb{I}(B)=\mathbb{I}_{\mathrm{p}}^{\infty}(B)$, that is, any ideal is generated by the projections that it contains. Then an ideal $I\mathrel{\triangleleft}B$ belongs to $\mathbb{I}_{\mathrm{p}}(B)$ if and only if $\widehat{I}$ is quasi-compact by Lemma 4.4. Any $\textup{C}^{*}$-algebra of real rank 00 has the ideal property by [Brown-Pedersen:C*-algebras_of_RR0]*Corollary 2.8 and Proposition 2.9.
Remark 4.6.
If $B$ is separable and purely infinite, then $I\mathrel{\triangleleft}B$ belongs to $\mathbb{I}_{\mathrm{p}}(B)$ if and only if $\widehat{I}$ is quasi-compact (see [Pasnicu-Rordam:Purely_infinite_rr0]*Proposition 2.7).
Definition 4.7 ([Ara-Moreno-Pardo:Nonstable_K_for_Graph_Algs, Brown-Pedersen:Non-stable]).
A $\textup{C}^{*}$-algebra $B$ has stable weak cancellation if projections $p,q$ in $\mathbb{M}_{\infty}(B)$ with $\operatorname{supp}p=\operatorname{supp}q$ and $[p]=[q]$ in $\textup{K}_{0}(\operatorname{supp}p)$ are equivalent.
The equality $[p]=[q]$ in $\textup{K}_{0}(\operatorname{supp}p)$ means that there is a projection $r\in\mathbb{M}_{\infty}(\operatorname{supp}p)$ so that $p\oplus r$ and $q\oplus r$ are equivalent, briefly, $p\oplus r\sim q\oplus r$. So stable weak cancellation holds if and only if $p\sim q$ follows if $p\oplus r\sim q\oplus r$ for a projection $r\in\mathbb{M}_{\infty}(\operatorname{supp}p\cap\operatorname{supp}q)$: this can only happen if $\operatorname{supp}p=\operatorname{supp}q$.
If $B$ is simple, then either $p=0$ or $\operatorname{supp}p=B$. Thus stable weak cancellation for simple $B$ means cancellation for all nonzero projections. It allows that there is a full (equivalently, nonzero) projection $0^{\mathrm{f}}\in\mathcal{P}(B)$ representing 00 in $\textup{K}_{0}(B)$. Then $\mathcal{P}(B)$ does not have cancellation because $0^{\mathrm{f}}+0^{\mathrm{f}}=0+0^{\mathrm{f}}$ in $\mathcal{P}(B)$. Thus stable weak cancellation is strictly weaker than cancellation in the monoid $\mathcal{P}(B)$.
Graph $\textup{C}^{*}$-algebras have stable weak cancellation by [Ara-Moreno-Pardo:Nonstable_K_for_Graph_Algs]*Corollary 7.2.
Theorem 4.8.
Purely infinite $\textup{C}^{*}$-algebras have stable weak cancellation.
Proof.
This follows from [Rordam:Classification_survey]*Theorem 4.1.4. ∎
Our next goal is to prove the following: if $B$ has stable weak cancellation, then the monoid $\mathcal{P}(B)$ contains the same information as the diagram of monoids $\textup{K}_{0}^{+}(I)$ for $I\in\mathbb{I}_{\mathrm{p}}(B)$ with the canonical maps $\textup{K}_{0}^{+}(I)\to\textup{K}_{0}^{+}(J)$ for $I,J\in\mathbb{I}_{\mathrm{p}}(B)$ with $I\subseteq J$. We first make the statement more precise.
Definition 4.9.
Let $(L,\leq)$ be a partially ordered set. A diagram of (semi)groups over it consists of (semi)groups $X_{W}$ for $W\in L$ and homomorphisms $\iota_{W_{2},W_{1}}\colon X_{W_{1}}\to X_{W_{2}}$ for $W_{1}\leq W_{2}$ in $L$, such that $\iota_{W,W}=\textup{id}_{X_{W}}$ for all $W\in L$ and $\iota_{W_{3},W_{2}}\circ\iota_{W_{2},W_{1}}=\iota_{W_{3},W_{1}}$ if $W_{1}\leq W_{2}\leq W_{3}$ in $L$.
Given diagrams $X=(X_{W},\varphi_{W_{2},W_{1}})$ and $Y=(Y_{W},\psi_{W_{2},W_{1}})$ over $L$, we let $\operatorname{Hom}_{L}(X,Y)$ be the (semi)group of natural transformations of diagrams; its elements are families of homomorphisms $X_{W}\to Y_{W}$ for $W\in L$ that intertwine the maps $\varphi_{W_{2},W_{1}}$ and $\psi_{W_{2},W_{1}}$ for $W_{1}\subseteq W_{2}$ in $L$.
Let $(L_{1},\leq)$ and $(L_{2},\leq)$ be partially ordered sets and $\psi\colon L_{1}\to L_{2}$ an order-preserving map. Let $(X_{W},\varphi_{W_{2},W_{1}})_{W\in L_{2}}$ be a diagram of (semi)groups over $L_{2}$. Then $\psi^{*}((X_{W},\varphi_{W_{2},W_{1}})_{W\in L_{2}})$ is the diagram of (semi)groups over $L_{1}$ consisting of the (semi)groups $X_{\psi(W)}$ for $W\in L_{1}$ and the homomorphisms $\varphi_{\psi(W_{2}),\psi(W_{1})}$ for $W_{1},W_{2}\in L_{1}$ with $W_{1}\leq W_{2}$.
Definition 4.10.
Let $B$ be a $\textup{C}^{*}$-algebra. Let $\textup{K}_{0}^{+}(B,\mathbb{I}_{\mathrm{p}})$ be the diagram over $\mathbb{I}_{\mathrm{p}}(B)$ consisting of the Abelian monoids $\textup{K}_{0}^{+}(I)$ for $I\in\mathbb{I}_{\mathrm{p}}(B)$ and the homomorphisms $\textup{K}_{0}^{+}(I)\to\textup{K}_{0}^{+}(J)$ induced by the inclusion maps $I\to J$ for $I,J\in\mathbb{I}_{\mathrm{p}}(B)$ with $I\subseteq J$. We define $\textup{K}_{0}(B,\mathbb{I}_{\mathrm{p}})$ and $\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$ similarly, replacing $\textup{K}_{0}^{+}$ by $\textup{K}_{0}$ and $\textup{K}_{1}$.
We recall the definition of an order unit:
Definition 4.11.
An order unit of a commutative semigroup $M$ (with its natural preorder $\leq$) is an element $u$ of $M$ such that for any element $x$ of $M$, there is a positive integer $n$ with $x\leq n\cdot u$. A homomorphism $\varphi\colon M\to M^{\prime}$ preserves order units if $\varphi(u)$ for an order unit $u$ in $M$ is an order unit in $M^{\prime}$. (If $\varphi(u)$ is an order unit for some order unit $u$ in $M$, then this happens for all order units $u$ in $M$.)
Let $A$ and $B$ be $\textup{C}^{*}$-algebras and let $\psi\colon\mathbb{I}_{\mathrm{p}}(A)\to\mathbb{I}_{\mathrm{p}}(B)$ be monotone. Then the pull-back diagram $\psi^{*}\bigl{(}\textup{K}_{0}^{+}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$ consists of the monoids $\textup{K}_{0}^{+}\bigl{(}\psi(I)\bigr{)}$ for $I\in\mathbb{I}_{\mathrm{p}}(A)$ with the canonical maps $\textup{K}_{0}^{+}\bigl{(}\psi(I_{1})\bigr{)}\to\textup{K}_{0}^{+}\bigl{(}\psi(I_{2})\bigr{)}$ for $I_{1}\subseteq I_{2}$.
Theorem 4.12.
Let $A$ and $B$ be $\textup{C}^{*}$-algebras. Let $B$ have stable weak cancellation. Then there is a natural bijection between monoid homomorphisms $f\colon\mathcal{P}(A)\to\mathcal{P}(B)$ and pairs $(\psi,(\varphi_{I})_{I\in\mathbb{I}_{\mathrm{p}}(A)})$, where $\psi\colon\mathbb{I}_{\mathrm{p}}(A)\to\mathbb{I}_{\mathrm{p}}(B)$ is monotone and
$\varphi=(\varphi_{I})_{I\in\mathbb{I}_{\mathrm{p}}(A)}\colon\textup{K}_{0}^{+}(A,\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{0}^{+}(B,\mathbb{I}_{\mathrm{p}})$
is a morphism of monoid diagrams such that each $\varphi_{I}$ preserves order units. The map $\psi$ in such a pair is automatically additive, that is, $\psi(I_{1}+I_{2})=\psi(I_{1})+\psi(I_{2})$ for all $I_{1},I_{2}\in\mathbb{I}_{\mathrm{p}}(A)$.
A monoid isomorphism clearly preserves order units. Hence Theorem 4.12 implies that $\mathcal{P}(A)\cong\mathcal{P}(B)$ if and only if there are an isomorphism of partially ordered sets $\psi\colon\mathbb{I}_{\mathrm{p}}(A)\xrightarrow{\sim}\mathbb{I}_{\mathrm{p}}(B)$ and an isomorphism $\textup{K}_{0}^{+}(A,\mathbb{I}_{\mathrm{p}})\cong\psi^{*}\bigl{(}\textup{K}_{0}^{+}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$ as diagrams of monoids over $\mathbb{I}_{\mathrm{p}}(A)$. This is what was meant by saying that $\mathcal{P}(A)$ and the diagram $\textup{K}_{0}^{+}(I)_{I\in\mathbb{I}_{\mathrm{p}}(A)}$ contain equivalent information.
We prepare the proof of Theorem 4.12 by two lemmas.
Lemma 4.13.
Let $B$ be a $\textup{C}^{*}$-algebra, $I\in\mathbb{I}_{\mathrm{p}}(B)$, and $p\in\mathcal{P}(I)$. Then $\operatorname{supp}p=I$ if and only if $p$ is an order unit in $\mathcal{P}(I)$. When $B$ has stable weak cancellation, this happens if and only if the image of $p$ in $\textup{K}_{0}^{+}(I)$ is an order unit.
Proof.
Since $I\in\mathbb{I}_{\mathrm{p}}(B)$, there is $q\in\mathcal{P}(B)$ with $I=\operatorname{supp}q$. We fix such a $q$ throughout the proof.
Assume that $p$ is an order unit in $\mathcal{P}(I)$. Then $q\leq n\cdot p$ for some $n\in\mathbb{N}_{>0}$ and hence $I=\operatorname{supp}q\subseteq n\cdot\operatorname{supp}p=\operatorname{supp}p$. Then $I=\operatorname{supp}p$. Conversely, let $\operatorname{supp}p=I$. Let $x\in\mathcal{P}(I)$. We must show that $x\leq n\cdot p$ for some $n\in\mathbb{N}_{>0}$. Equivalently, $x\cdot I^{\infty}$ is a direct summand in $(p\cdot I^{\infty})^{n}$ as a Hilbert $I$-module.
Let :$\mathcal{H}\mathrel{\vcentcolon=}p\cdot I^{\infty}$ as a right Hilbert $I$-module. Then $\mathbb{K}(\mathcal{H})\cong p(I\otimes\mathbb{K})p$. Let $\mathcal{H}^{*}$ be the dual Hilbert $I,\mathbb{K}(\mathcal{H})$-bimodule. Then $\mathcal{H}^{*}\otimes_{\mathbb{K}(\mathcal{H})}\mathcal{H}\cong I$ and $\mathcal{H}\otimes_{I}\mathcal{H}^{*}\cong\mathbb{K}(\mathcal{H})$. So $x\cdot I^{\infty}$ is a direct summand in $(p\cdot I^{\infty})^{n}$ if and only if $x\cdot I^{\infty}\otimes_{I}\mathcal{H}^{*}$ is a direct summand in $(p\cdot I^{\infty})^{n}\otimes_{I}\mathcal{H}^{*}=\mathbb{K}(\mathcal{H})^{n}$ as a Hilbert $\mathbb{K}(\mathcal{H})$-module. The $\textup{C}^{*}$-algebra $\mathbb{K}(x\cdot I^{\infty}\otimes_{I}\mathcal{H}^{*})$ is isomorphic to $\mathbb{K}(x\cdot I^{\infty})$ and hence still unital. So $x\cdot I^{\infty}\otimes_{I}\mathcal{H}^{*}$ is a proper Hilbert $\mathbb{K}(\mathcal{H})$-module. By Theorem 2.3, $x\cdot I^{\infty}\otimes_{I}\mathcal{H}^{*}$ is a direct summand in $(p\cdot I^{\infty})^{n}$ for some $n\in\mathbb{N}_{>0}$. This finishes the proof that $\operatorname{supp}p=I$ if and only if $p$ is an order unit in $\mathcal{P}(I)$.
Let $p$ be an order unit in $\mathcal{P}(I)$. If $x\in\textup{K}_{0}^{+}(I)$, then $x=[\bar{x}]$ for some $\bar{x}\in\mathcal{P}(I)$. By assumption, $\bar{x}\leq n\cdot p$ for some $n\in\mathbb{N}_{>0}$. Hence $x\leq n\cdot[p]$. Thus $[p]\in\textup{K}_{0}^{+}(I)$ is an order unit in $\textup{K}_{0}^{+}(I)$. Conversely, let $[p]\in\textup{K}_{0}^{+}(I)$ be an order unit. Let $x\in\mathcal{P}(I)$. Then $x\oplus q$ is full in $I$. There is $n\in\mathbb{N}_{>0}$ with $[x\oplus q]\leq n\cdot[p]$ in $\textup{K}_{0}^{+}(I)$, that is, $[x\oplus q]\oplus[y]=n\cdot[p]$ for some $y\in\mathcal{P}(I)$. Since $B$ has stable weak cancellation and both sides are full in $I$, this implies $x\oplus q\oplus y\sim n\cdot p$ in $\mathcal{P}(I)$. So $x\leq n\cdot p$. Thus $p$ is also an order unit in $\mathcal{P}(I)$. ∎
Definition 4.14.
For $p,q\in\mathcal{P}(B)$, we write $p\prec q$ if and only if $\operatorname{supp}p\subseteq\operatorname{supp}q$.
Lemma 4.15.
Let $p,q\in\mathcal{P}(B)$. Then $p\prec q$ if and only if $p\leq n\cdot q$ for some $n\in\mathbb{N}$.
Proof.
If $p\leq n\cdot q$, then $\operatorname{supp}p\subseteq\operatorname{supp}n\cdot q=\operatorname{supp}q$. Conversely, let $\operatorname{supp}p\subseteq\operatorname{supp}q$. By Lemma 4.13, $q$ is an order unit for $\mathcal{P}(\operatorname{supp}q)$. Hence $p\leq n\cdot q$ for some $n\in\mathbb{N}$. ∎
Proof of Theorem 4.12.
Let $f\colon\mathcal{P}(A)\to\mathcal{P}(B)$ be a homomorphism. The map $\operatorname{supp}\colon\mathcal{P}(A)\to(\mathbb{I}_{\mathrm{p}}(A),+)$ is a surjective monoid homomorphism. Hence $\mathbb{I}_{\mathrm{p}}(A)$ is isomorphic to the quotient of $\mathcal{P}(A)$ by the congruence relation defined by $\operatorname{supp}p=\operatorname{supp}q$. Lemma 4.15 describes this congruence relation and also the partial order defined by $\operatorname{supp}p\subseteq\operatorname{supp}q$ in a purely algebraic way. So $\operatorname{supp}p\subseteq\operatorname{supp}q$ implies $\operatorname{supp}f(p)\subseteq\operatorname{supp}f(q)$ and similarly with equality instead of inclusion. As a result, $f$ induces a well defined map $\psi\colon\mathbb{I}_{\mathrm{p}}(A)\to\mathbb{I}_{\mathrm{p}}(B)$ that is both additive and monotone.
Let $I\in\mathbb{I}_{\mathrm{p}}(A)$. Then there is $q\in\mathcal{P}(A)$ with $I=\operatorname{supp}q$. Then $\psi(I)=\operatorname{supp}f(q)$. Lemma 4.15 shows that $p\in\mathcal{P}(A)$ satisfies $p\in\mathcal{P}(I)$ if and only if there is $n\in\mathbb{N}$ with $p\leq n\cdot q$. Hence $f$ maps $\mathcal{P}(I)$ to $\mathcal{P}(\psi(I)))$. This induces a group homomorphism $\textup{K}_{0}(I)\to\textup{K}_{0}(\psi(I))$, which restricts to a monoid homomorphism $\varphi_{I}\colon\textup{K}_{0}^{+}(I)\to\textup{K}_{0}^{+}(\psi(I))$. It maps the order unit $q$ in $\mathcal{P}(I)$ to $f(q)$, which is an order unit in $\mathcal{P}(\psi(I))$ by Lemma 4.13 because $\psi(I)=\operatorname{supp}f(q)$. These homomorphisms for $I\in\mathbb{I}_{\mathrm{p}}(A)$ form a morphism of diagrams over $\mathbb{I}_{\mathrm{p}}(A)$ that preserves order units.
Conversely, let $\psi\colon\mathbb{I}_{\mathrm{p}}(A)\to\mathbb{I}_{\mathrm{p}}(B)$ be a monotone map and let $\varphi_{I}\colon\textup{K}_{0}^{+}(I)\to\textup{K}_{0}^{+}(\psi(I))$ for $I\in\mathbb{I}_{\mathrm{p}}(A)$ be monoid homomorphisms that preserve order units and form a morphism of diagrams. Assume that $B$ has stable weak cancellation. Let $p\in\mathcal{P}(A)$. Let $I\mathrel{\vcentcolon=}\operatorname{supp}p$. Choose $q\in\mathcal{P}(\psi(I))$ representing $\varphi_{I}[p]$ in $\textup{K}_{0}^{+}(\psi(I))$. Since $\varphi_{I}$ preserves order units, $\operatorname{supp}q=\psi(I)$. Since $B$ has stable weak cancellation, $q$ is unique as an element of $\mathcal{P}(B)$. So $f(p)\mathrel{\vcentcolon=}q$ well-defines a map $f\colon\mathcal{P}(A)\to\mathcal{P}(B)$.
Let $p_{1},p_{2}\in\mathcal{P}(A)$. We claim that $f(p_{1}+p_{2})=f(p_{1})+f(p_{2})$. Let $p_{3}\mathrel{\vcentcolon=}p_{1}+p_{2}$. For $j=1,2,3$, define $I_{j}\mathrel{\vcentcolon=}\operatorname{supp}p_{j}$. By Lemma 4.13, $[p_{j}]\in\textup{K}_{0}^{+}(I_{j})$ is an order unit. Therefore, $\varphi_{I_{j}}[p_{j}]\in\textup{K}_{0}^{+}(\psi(I_{j}))$ is an order unit. Since $B$ has stable weak cancellation, $q_{j}\mathrel{\vcentcolon=}f(p_{j})\in\mathcal{P}(B)$ is the unique element with support ideal $\psi(I_{j})$ and $\varphi_{I_{j}}[p_{j}]=[q_{j}]$ in $\textup{K}_{0}^{+}(\psi(I_{j}))$. The relation $p_{1}+p_{2}=p_{3}$ holds in $\textup{K}_{0}^{+}(I_{3})$ when we tacitly view $p_{1},p_{2},p_{3}$ as elements of $\mathcal{P}(I_{3})$. Since $\varphi_{I_{3}}$ is additive, this implies the relation $[q_{1}+q_{2}]=[q_{1}]+[q_{2}]=[q_{3}]$ in $\textup{K}_{0}^{+}(\psi(I_{3}))$. Since the maps $\varphi_{I}$ preserve order units, $[q_{1}+q_{2}]=[q_{3}]$ is an order unit for $\textup{K}_{0}^{+}(\psi(I_{3}))$. Then Lemma 4.13 implies that $\psi(I_{3})$ is the support ideal of $q_{3}$ and of $q_{1}+q_{2}$. The support of $q_{1}+q_{2}$ is contained in $\psi(I_{1})+\psi(I_{2})$. So $\psi(I_{3})\subseteq\psi(I_{1})+\psi(I_{2})$. Since $\psi$ is monotone by assumption, it follows that $\psi$ must be additive. And stable weak cancellation for $B$ shows that $q_{1}+q_{2}=q_{3}$ holds in $\mathcal{P}(B)$. That is, $f\colon\mathcal{P}(A)\to\mathcal{P}(B)$ is a homomorphism. It is easy to see that the two constructions above are inverse to each other, giving the asserted bijection. ∎
Corollary 4.16.
Let $A$ and $B$ be $\textup{C}^{*}$-algebras. Assume that $B$ is purely infinite. There is a natural bijection between homomorphisms $f\colon\mathcal{P}(A)\to\mathcal{P}(B)$ and pairs $(\psi,(\varphi_{I})_{I\in\mathbb{I}_{\mathrm{p}}(A)})$ with a monotone map $\psi\colon\mathbb{I}_{\mathrm{p}}(A)\to\mathbb{I}_{\mathrm{p}}(B)$ and a morphism of diagrams of Abelian groups $\varphi_{I}\colon\textup{K}_{0}(A,\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{0}(B,\mathbb{I}_{\mathrm{p}})$.
Proof.
Since any $I\in\mathbb{I}_{\mathrm{p}}(B)$ is also purely infinite, any class in $\textup{K}_{0}(I)$ is represented by a projection in $I$ (see [Rordam:Classification_survey]*Theorem 4.1.4). Thus $\textup{K}_{0}^{+}(B,\mathbb{I}_{\mathrm{p}})=\textup{K}_{0}(B,\mathbb{I}_{\mathrm{p}})$, which is a diagram of Abelian groups. Any element of an Abelian group is an order unit. Theorem 4.8 says that $B$ has stable weak cancellation. So Theorem 4.12 applies here and gives the remaining statements with $\textup{K}_{0}^{+}$ instead of $\textup{K}_{0}$. If $I\in\mathbb{I}_{\mathrm{p}}(A)$, then $\textup{K}_{0}(I)$ is isomorphic to the Grothendieck group of $\textup{K}_{0}^{+}(I)$. Hence there is a bijection between homomorphisms $\textup{K}_{0}^{+}(I)\to\textup{K}_{0}^{+}(J)$ and $\textup{K}_{0}(I)\to\textup{K}_{0}(J)$ for $I\in\mathbb{I}_{\mathrm{p}}(A)$ and $J\in\mathbb{I}_{\mathrm{p}}(B)$. ∎
Remark 4.17.
[…]
5. Some homological algebra with K-theory diagrams
We first recall well known results on the ideal structure of graph $\textup{C}^{*}$-algebras and about their K-theory diagrams $\textup{K}_{0}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})$, $\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})$ and $\textup{K}_{0}^{+}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})$.
Definition 5.1.
A subset $H\subseteq V$ is hereditary if for any edge $e\in E$, $r(e)\in H$ implies $s(e)\in H$. It is saturated if any $v\in V_{\mathrm{reg}}$ with $s(e)\in H$ for all edges $e$ with $r(e)=v$ satisfies $v\in H$.
The intersection of any set of hereditary and saturated subsets is again hereditary and saturated. The union may fail to be hereditary and saturated, but then there is a smallest hereditary and saturated subset containing it. Thus the hereditary and saturated subsets of $V$ form a complete lattice, which we denote by $\mathbb{I}(G)$.
Definition 5.2.
For a subset $V_{2}\subseteq V$, let $\langle V_{2}\rangle\subseteq V$ be the smallest hereditary and saturated subset containing $V_{2}$. If $V_{2}=\{v\}$, we briefly denote this by $\langle v\rangle$.
There is an order-isomorphism between $\mathbb{I}(G)$ and the lattice $\mathbb{I}^{\mathbb{T}}(\textup{C}^{*}(G))$ of gauge-invariant ideals in $\textup{C}^{*}(G)$ because $G$ is row-finite (see [Bates-Pask-Raeburn-Szymanski:Row_finite]*Thereom 4.1 or [Bates-Hong-Raeburn-Szymanski:Ideal_structure]*Theorem 3.6). It maps a hereditary and saturated subset $H\subseteq V$ to the ideal generated by the projections $p_{v}$ for all $v\in H$. Its inverse maps a gauge-invariant ideal $I$ to the set of $v\in V$ with $p_{v}\in I$. We write $\textup{C}^{*}(G)_{H}$ for the ideal corresponding to $H\in\mathbb{I}(G)$. It is Morita–Rieffel equivalent to the graph $\textup{C}^{*}$-algebra of the restriction $G|_{H}$ of $G$ to $H$.
Write $q\in\mathcal{P}(\textup{C}^{*}(G))$ as $\sum_{v\in V}c_{v}p_{v}$ as in Theorem 3.1. Then $\operatorname{supp}q$ is equal to the ideal generated by the projections $p_{v}$ with $c_{v}\neq 0$. An ideal of this form is automatically gauge-invariant. The hereditary, saturated subset of $V$ that corresponds to $\operatorname{supp}q$ is the smallest one that contains all $v$ with $c_{v}\neq 0$. It is called the (hereditary, saturated) support of $c$. The following theorem summarises the findings:
Theorem 5.3.
Let $G$ be a graph. An ideal of $\textup{C}^{*}(G)$ belongs to $\mathbb{I}_{\mathrm{p}}^{\infty}(\textup{C}^{*}(G))$ if and only if it is gauge-invariant, and $\mathbb{I}_{\mathrm{p}}^{\infty}(\textup{C}^{*}(G))\cong\mathbb{I}(G)$ as complete lattices. This isomorphism maps the subset $\mathbb{I}_{\mathrm{p}}(\textup{C}^{*}(G))$ onto the partially ordered set of finitely generated hereditary, saturated subsets of $V$; we denote this by $\mathbb{I}_{\mathrm{p}}(G)$. The support map $\operatorname{supp}\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathbb{I}_{\mathrm{p}}(\textup{C}^{*}(G))\cong\mathbb{I}_{\mathrm{p}}(G)$ maps the class of the projection $\sum c_{v}p_{v}$ for $c\in\mathbb{N}[V]$ to the hereditary, saturated support of $c$.
The free Abelian groups $\mathbb{Z}[W]$ for $W\in\mathbb{I}_{\mathrm{p}}(G)$ form a diagram of groups over $\mathbb{I}_{\mathrm{p}}(G)$, which we denote by $\mathbb{Z}[V,\mathbb{I}_{\mathrm{p}}]$. Similarly, let $W_{\mathrm{reg}}\mathrel{\vcentcolon=}W\cap V_{\mathrm{reg}}$ for $W\in\mathbb{I}_{\mathrm{p}}(G)$. The free Abelian groups $\mathbb{Z}[W_{\mathrm{reg}}]$ for $W\in\mathbb{I}_{\mathrm{p}}(G)$ form a diagram of groups over $\mathbb{I}_{\mathrm{p}}(G)$ denoted by $\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}]$. We define a morphism of diagrams $M_{G}^{\mathbb{I}_{\mathrm{p}}}\colon\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}]\to\mathbb{Z}[V,\mathbb{I}_{\mathrm{p}}]$ to consist of the maps
$M_{G}^{W}\colon\mathbb{Z}[W_{\mathrm{reg}}]\to\mathbb{Z}[W],\qquad M_{G}^{W}\bigl{(}(c_{w})_{w\in W}\bigr{)}(v)\mathrel{\vcentcolon=}\sum_{\{e\in E\,{:}\,\mathopen{}s(e)=v\}}c_{r(e)}.$
These maps form a diagram morphism because if $W\subseteq V$ is hereditary, then $M_{G}^{W}=M_{G}^{V}|_{\mathbb{Z}[W]}$. Roughly speaking, $M_{G}^{\mathbb{I}_{\mathrm{p}}}$ multiplies with the adjacency matrix of $G$, that is, with the matrix whose $v,w$-entry is the number of edges $v\leftarrow w$ in $G$.
The results on the K-theory of graph $\textup{C}^{*}$-algebras provide natural exact sequences
(5.4) $0\to\textup{K}_{1}(\textup{C}^{*}(G|_{W}))\xrightarrow{\kappa|_{W}}\mathbb{Z}[W_{\mathrm{reg}}]\xrightarrow{\textup{id}-M_{G}^{W}}\mathbb{Z}[W]\xrightarrow{\pi|_{W}}\textup{K}_{0}(\textup{C}^{*}(G|_{W}))\to 0$
for all $W\in\mathbb{I}_{\mathrm{p}}^{\infty}(G)$ (see [Raeburn-Szymanski:CK_algs_of_inf_graphs_and_matrices]*Theorem 3.2). These are compatible with the diagram structure over $\mathbb{I}_{\mathrm{p}}(G)$, so that (5.4) gives an exact sequence of diagrams
[…]
In addition, $\textup{K}_{0}^{+}(\textup{C}^{*}(G|_{W}))=\pi|_{W}(\mathbb{N}[W])\subseteq\textup{K}_{0}(\textup{C}^{*}(G|_{W}))$ by Theorem 3.1.
Diagrams of Abelian groups over $\mathbb{I}_{\mathrm{p}}(G)$ form an Abelian category. This allows us to do homological algebra with them. We shall need the extension groups in this category, which we denote by $\operatorname{Ext}^{n}_{\mathbb{I}_{\mathrm{p}}(G)}$ for $n\in\mathbb{N}$. The case $n=0$ gives $\operatorname{Ext}^{0}=\operatorname{Hom}$.
Proposition 5.6.
The diagrams $\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}]$ and $\mathbb{Z}[V,\mathbb{I}_{\mathrm{p}}]$ are projective. So (5.5) is the beginning of a projective resolution of $\textup{K}_{0}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})$.
Proof.
For $v\in V$, let $\langle v\rangle\subseteq V$ be the hereditary and saturated subset generated by $v$. So $\langle v\rangle\in\mathbb{I}_{\mathrm{p}}(G)$. Let $(X_{W},\varphi_{W_{2},W_{1}})_{W\in\mathbb{I}_{\mathrm{p}}(G)}$ be a diagram of groups over $\mathbb{I}_{\mathrm{p}}(G)$. A family $x_{v}\in X_{\langle v\rangle}$ for $v\in V$ defines a diagram morphism $\xi\colon\mathbb{Z}[V,\mathbb{I}_{\mathrm{p}}]\to(X_{W})_{W\in\mathbb{I}_{\mathrm{p}}(G)}$, which consists of the maps
$\xi_{W}\colon\mathbb{Z}[W]\to X_{W},\qquad\sum_{v\in W}c_{v}\delta_{v}\mapsto\sum_{v\in W}c_{v}\varphi_{W,\langle v\rangle}(x_{v}).$
Any diagram morphism $\xi\colon\mathbb{Z}[V,\mathbb{I}_{\mathrm{p}}]\to(X_{W})_{W\in\mathbb{I}_{\mathrm{p}}(G)}$ is of this form for a unique family of $x_{v}\in X_{\langle v\rangle}$ for $v\in V$, namely, $x_{v}\mathrel{\vcentcolon=}\xi_{\langle v\rangle}(\delta_{v})$. So
$\operatorname{Hom}_{\mathbb{I}_{\mathrm{p}}(G)}\bigl{(}\mathbb{Z}[V,\mathbb{I}_{\mathrm{p}}],(X_{W})\bigr{)}\cong\prod_{v\in W}X_{\langle v\rangle}.$
This functor is exact. That is, $\mathbb{Z}[V,\mathbb{I}_{\mathrm{p}}]$ is projective. A similar argument works for $\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}]$ or, more generally, for the diagram $W\mapsto\mathbb{Z}[W\cap X,\mathbb{I}_{\mathrm{p}}]$ for any subset $X$ of $V$. ∎
The resolution in (5.5) allows us to compute $\operatorname{Ext}^{n}_{\mathbb{I}_{\mathrm{p}}(G)}(\textup{K}_{0}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}}),Y)$ for any diagram of groups $Y$ over $\mathbb{I}_{\mathrm{p}}(G)$ and $n=0,1,2$ as the cohomology groups of the cochain complex
(5.7) $0\to\operatorname{Hom}_{\mathbb{I}_{\mathrm{p}}(G)}(\mathbb{Z}[V,\mathbb{I}_{\mathrm{p}}],Y)\xrightarrow{(\textup{id}-M_{G}^{\mathbb{I}_{\mathrm{p}}})^{*}}\operatorname{Hom}_{\mathbb{I}_{\mathrm{p}}(G)}(\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}],Y)\\ \xrightarrow{\kappa^{*}}\operatorname{Hom}_{\mathbb{I}_{\mathrm{p}}(G)}(\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}}),Y)\to 0.$
This works for $n=0,1,2$, regardless whether or not $\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})$ is projective.
We now compare the resolution above to the constructions in [Bentmann-Meyer:More_general]. That article only considers graph $\textup{C}^{*}$-algebras with finitely many ideals and, without loss of generality, assumes all graphs to be regular. While each ideal in $\textup{C}^{*}(G)$ is generated by a projection in that case, only some ideals are used in [Bentmann-Meyer:More_general]. Call an ideal $I\mathrel{\triangleleft}B$ irreducible if $I=\overline{\sum_{\alpha\in S}J_{\alpha}}$ for a family of ideals $J_{\alpha}$ implies $J_{\alpha}=I$ for some $\alpha\in S$. The irreducible ideals form a partially ordered set. The invariant that is used in [Bentmann-Meyer:More_general] is the diagram consisting of the $\mathbb{Z}/2$-graded Abelian groups $\textup{K}_{*}(I)$ for all irreducible ideals $I\mathrel{\triangleleft}B$. If $B$ has only finitely many ideals, then for each primitive ideal $\mathfrak{p}\mathrel{\triangleleft}B$ there is a minimal open subset of the primitive ideal space that contains $\mathfrak{p}$. The ideals corresponding to these minimal open subsets are exactly the irreducible ideals.
Remark 5.8.
The Cantor set is the primitive ideal space of a commutative AF-algebra and hence of a graph $\textup{C}^{*}$-algebra. All open subsets of the Cantor set are reducible. So restricting to irreducible ideals is not reasonable for $\textup{C}^{*}$-algebras with infinitely many ideals, not even if they are graph $\textup{C}^{*}$-algebras.
For a graph $\textup{C}^{*}$-algebra with finitely many ideals, let $\mathbb{I}_{\mathrm{s}}(G)\subseteq\mathbb{I}_{\mathrm{p}}(G)=\mathbb{I}(G)$ be the subset corresponding to the irreducible ideals in $\textup{C}^{*}(G)$. We may restrict the exact sequence of diagrams (5.5) to the subset $\mathbb{I}_{\mathrm{s}}(G)\subseteq\mathbb{I}_{\mathrm{p}}(G)$. The restricted diagrams $\mathbb{Z}[V,\mathbb{I}_{\mathrm{s}}]$ and $\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{s}})$ are shown in [Bentmann-Meyer:More_general] to be projective objects in the Abelian category of diagrams of Abelian groups over $\mathbb{I}_{\mathrm{s}}(G)$. Hence the groups $\operatorname{Ext}^{n}_{\mathbb{I}_{\mathrm{s}}(G)}\bigl{(}\textup{K}_{0}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{s}}),\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{s}})\bigr{)}$ in this Abelian category of diagrams vanish for $n\geq 3$ and are computed exactly as in (5.7) with $\mathbb{I}_{\mathrm{s}}$ instead of $\mathbb{I}_{\mathrm{p}}$ if $n=0,1,2$.
The invariant over $\mathbb{I}_{\mathrm{s}}$ clearly contains less information than the invariant over $\mathbb{I}_{\mathrm{p}}$. This does not mean, however, that lifting a map on the invariant over $\mathbb{I}_{\mathrm{p}}$ is easier than lifting a map on the invariant over $\mathbb{I}_{\mathrm{s}}$: a given extension to $\mathbb{I}_{\mathrm{p}}$ may put constraints on the desired $\textup{C}^{*}$-correspondence.
Remark 5.9.
Let $G$ be a regular graph with only finitely many hereditary saturated subsets. Then any irreducible hereditary saturated subset is generated by a single vertex $v\in V$. The converse does not hold, however, because the union of two hereditary and saturated subsets of $V$ need not remain hereditary. For instance, let $G$ be the graph
312.
The hereditary and saturated subsets for $G$ are $\emptyset$, $\{1\}$, $\{2\}$, and $\{1,2,3\}$. The last one is singly generated by $3$, but not irreducible.
Remark 5.10.
The machinery of homological algebra in triangulated categories does not apply to the invariant $\textup{K}_{*}(B,\mathbb{I}_{\mathrm{p}})$, not even if the ideal lattice of $B$ is finite. This remark explains why. First we make our invariant more precise. Let $X$ be a locally quasi-compact topological space and let $(L,\leq)$ be the partially ordered set of quasi-compact open subsets of $X$. Let $B$ be a $\textup{C}^{*}$-algebra over $X$. If $\operatorname{Prim}(B)\cong X$ and $B$ has the ideal property, then $L=\mathbb{I}_{\mathrm{p}}(B)$. Diagrams of $\mathbb{Z}/2$-graded countable Abelian groups over $(L,\leq)$ form a stable Abelian category, and mapping $B$ to the diagram $\textup{K}_{*}(B,L)\mathrel{\vcentcolon=}\bigl{(}\textup{K}_{*}(I)\bigr{)}_{I\in L}$ is a stable homological functor into this category. This invariant is, however, not “universal,” so that the homological algebra machinery developed in [Meyer-Nest:Homology_in_KK, Meyer:Homology_in_KK_II, Meyer-Nest:Bootstrap, Meyer-Nest:Filtrated_K] does not apply directly to it. Let $I_{1},I_{2}\in L$ and assume $I_{1}\cap I_{2}\in L$. We have $I_{1}+I_{2}\in L$ anyway. The boundary map $\partial$ in the Mayer–Vietoris sequence
$\dotsb\to\textup{K}_{*}(I_{1}\cap I_{2})\to\textup{K}_{*}(I_{1})\oplus\textup{K}_{*}(I_{2})\to\textup{K}_{*}(I_{1}+I_{2})\xrightarrow{\partial}\textup{K}_{*-1}(I_{1}\cap I_{2})\to\dotsb$
is a natural transformation from $\textup{K}_{*}(I_{1}+I_{2})$ to $\textup{K}_{*-1}(I_{1}\cap I_{2})$, which is not yet encoded in our Abelian category of diagrams. Hence our invariant cannot be universal.
Despite Remark 5.10, we shall work with the invariant $\textup{K}_{*}(B,\mathbb{I}_{\mathrm{p}})$ and use homological algebra in the category of diagrams over $\mathbb{I}_{\mathrm{p}}(G)$. This works quite well to describe correspondences from graph $\textup{C}^{*}$-algebras to $B$. In fact, we shall use $\mathcal{P}(B)$ instead of $\textup{K}_{0}^{+}(B,\mathbb{I}_{\mathrm{p}})$ most of the time. These invariants are equivalent by Theorem 4.12 if $B$ has stable weak cancellation. We do not need to assume stable weak cancellation when we work directly with $\mathcal{P}(B)$.
6. Adding $\textup{K}_{1}$
Theorem 3.2 only uses the invariant $\mathcal{P}(B)$, which is not fine enough for classification. Now we add information about $\textup{K}_{1}$. We develop a criterion when a pair of maps $\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$ and $\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$ is induced by a proper $\textup{C}^{*}(G),B$-correspondence. It uses a variant of the obstruction class in [Bentmann-Meyer:More_general].
Let $\mathcal{H}$ be a proper $\textup{C}^{*}(G),B$-correspondence. Describe $\mathcal{H}$ through proper Hilbert $B$-modules $\mathcal{H}_{v}$ for $v\in V$ and unitaries $U_{v}\colon\bigoplus_{e\in E^{v}}\mathcal{H}_{s(e)}\xrightarrow{\sim}\mathcal{H}_{v}$ for $v\in V_{\mathrm{reg}}$ as in Theorem 3.2. We are going to describe the homomorphisms
$\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B),\quad\mathbb{I}(\textup{C}^{*}(G))\to\mathbb{I}_{\mathrm{p}}(B),\quad\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$
induced by $\mathcal{H}$. By Theorem 3.1, the homomorphism $\mathcal{H}_{*}\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$, $[\mathcal{M}]\mapsto[\mathcal{M}\otimes_{\textup{C}^{*}(G)}\mathcal{H}]$, is determined by its values
$\mathcal{H}_{*}[p_{v}]=\mathcal{H}_{*}[p_{v}\cdot\textup{C}^{*}(G)]=[p_{v}\cdot\mathcal{H}]=[\mathcal{H}_{v}].$
The homomorphism $\mathcal{H}_{*}\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$ induces a map
$\psi\colon\mathbb{I}_{\mathrm{p}}(G)\cong\mathbb{I}_{\mathrm{p}}(\textup{C}^{*}(G))\to\mathbb{I}_{\mathrm{p}}(B)$
as in the proof of Theorem 4.12; this part of the proof works without stable weak cancellation. Namely, let $W\in\mathbb{I}_{\mathrm{p}}(G)$ be a finitely generated, hereditary, saturated subset of $V$. Let $F\subseteq W$ be any finite generating set for $W$. Then the projection $\bigoplus_{v\in F}p_{v}$ generates the ideal $\textup{C}^{*}(G)_{W}$ corresponding to $W$. Hence
$\psi(W)\mathrel{\vcentcolon=}\operatorname{supp}\varphi_{0}\Bigl{(}\sum_{v\in F}{}[p_{v}]\Bigr{)}=\bigvee_{v\in F}{}\operatorname{supp}\varphi_{0}[p_{v}]$
This ideal is equal to $\bigvee_{v\in W}{}\operatorname{supp}\varphi_{0}[p_{v}]$ and does not depend on the choice of $F$.
By definition, $\psi(W)$ is the ideal in $B$ that is generated by the inner product on the proper Hilbert $B$-module $\bigoplus_{v\in F}p_{v}\cdot\mathcal{H}=\bigoplus_{v\in F}\mathcal{H}_{v}$. Therefore, $\textup{C}^{*}(G)_{W}\cdot\mathcal{H}$ is a full Hilbert $\psi(W)$-module. As such, it is nondegenerate as a right $\psi(W)$-module, so that
(6.1) $\textup{C}^{*}(G)_{W}\cdot\mathcal{H}\subseteq\mathcal{H}\cdot\psi(W).$
We briefly call a $\textup{C}^{*}(G),B$-correspondence $\mathcal{H}$ satisfying (6.1) $\mathbb{I}_{\mathrm{p}}(G)$-covariant.
Equation (6.1) implies that $\mathcal{H}$ restricts to a proper $\textup{C}^{*}(G)_{W},\psi(W)$-correspondence and thus induces a grading-preserving group homomorphism
$\mathcal{H}_{*}|_{W}\colon\textup{K}_{*}(\textup{C}^{*}(G)_{W})\to\textup{K}_{*}(\psi(W)).$
These homomorphisms for $W\in\mathbb{I}_{\mathrm{p}}(G)$ form a morphism of diagrams
$\mathcal{H}_{*}\colon\textup{K}_{*}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{*}(B,\mathbb{I}_{\mathrm{p}}).$
We shall not use this on $\textup{K}_{0}$ in the following because it contains less information than the homomorphism $\mathcal{H}_{*}\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$ (compare Theorem 4.12). We now describe the map $\mathcal{H}_{*}|_{W}\colon\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$. Since the map $\kappa|_{W}$ in (5.4) is injective, we may identify $\textup{K}_{1}(\textup{C}^{*}(G)_{W})\cong\textup{K}_{1}(\textup{C}^{*}(G|_{W}))$ with the kernel of the map $\textup{id}-M_{G}^{W}\colon\mathbb{Z}[W_{\mathrm{reg}}]\to\mathbb{Z}[W]$. Let $x=\sum c_{v}\delta_{v}\in\mathbb{Z}[W_{\mathrm{reg}}]$ belong to this kernel. Write $x=\sum c_{v}^{+}\delta_{v}-\sum c_{v}^{-}\delta_{v}$ with $c_{v}^{\pm}\geq 0$, $c_{v}=c_{v}^{+}-c_{v}^{-}$ and $c_{v}^{+}=0$ or $c_{v}^{-}=0$ for all $v\in W_{\mathrm{reg}}$. Set $c_{v}=0$ for $v\in V_{\mathrm{reg}}\setminus W$. The equation $(\textup{id}-M_{G}^{W})(x)=0$ says that
(6.2) $\sum_{v\in V_{\mathrm{reg}}}c_{v}^{+}\delta_{v}+\sum_{e\in E}c_{r(e)}^{-}\delta_{s(e)}=\sum_{e\in E}c_{r(e)}^{+}\delta_{s(e)}+\sum_{v\in V_{\mathrm{reg}}}c_{v}^{-}\delta_{v}.$
Equivalently, the domains and codomains of the partial isometry
$u_{x}\mathrel{\vcentcolon=}\bigoplus_{v\in W_{\mathrm{reg}}}u_{v}^{\oplus c_{v}^{+}}\oplus\bigoplus_{v\in W_{\mathrm{reg}}}(u_{v}^{*})^{\oplus c_{v}^{-}}$
are equal up to a permutation of the summands. Let $u^{\prime}_{x}$ be the product of $u_{x}$ with the permutation matrix that makes it a unitary element in the corner of $\mathbb{M}_{\infty}(\textup{C}^{*}(G)_{W})$ associated to the projection
$p_{x}\mathrel{\vcentcolon=}\bigoplus_{v\in V_{\mathrm{reg}}}p_{v}^{\oplus c_{v}^{+}}+\bigoplus_{e\in E}p_{s(e)}^{\oplus c_{r(e)}^{-}}.$
Then $u^{\prime}_{x}$ has a class in $\textup{K}_{1}(\textup{C}^{*}(G)_{W})$. The exact sequence (5.5) is proved by showing that any class in $\textup{K}_{1}(\textup{C}^{*}(G)_{W})$ is of this form for a unique $x\in\ker(\textup{id}-M_{G}^{W})$. Let $P_{v}\in\mathbb{K}(\mathcal{H})$ be the projection onto $\mathcal{H}_{v}$ and let $U_{v}$ be as in Theorem 3.2. Define
$P_{x}\mathrel{\vcentcolon=}\bigoplus_{v\in V_{\mathrm{reg}}}P_{v}^{\oplus c_{v}^{+}}\oplus\bigoplus_{e\in E}P_{s(e)}^{\oplus c_{r(e)}^{-}},$
and define a unitary $U^{\prime}_{x}$ in $P_{x}\mathbb{K}(\mathcal{H}^{\infty})P_{x}$ in the same way as $u^{\prime}_{x}$. This is how $u^{\prime}_{x}\in p_{x}\mathbb{M}_{\infty}(\textup{C}^{*}(G))p_{x}$ acts on $P_{x}\mathcal{H}^{\infty}$, and the latter is a Hilbert $\psi(W)$-module. The map $\mathcal{H}_{*}|_{W}\colon\textup{K}_{1}(\textup{C}^{*}(G)_{W})\to\textup{K}_{1}(\psi(W))$ maps $x$ to $[U^{\prime}_{x}]$.
Now let $\varphi_{0}\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$ be a monoid homomorphism. Let
$\psi\colon\mathbb{I}_{\mathrm{p}}(G)\cong\mathbb{I}_{\mathrm{p}}(\textup{C}^{*}(G))\to\mathbb{I}_{\mathrm{p}}(B)$
be induced by $\varphi_{0}$. And let $\varphi_{1}\colon\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$ be a morphism of diagrams over $\mathbb{I}_{\mathrm{p}}(G)$. When is there a proper $\textup{C}^{*}(G),B$-correspondence $\mathcal{H}$ that induces the maps $\varphi_{0}$ and $\varphi_{1}$? Our criterion needs a variant of the obstruction class in [Bentmann-Meyer:More_general]. Our variant lives in the $\operatorname{Ext}^{2}$-group
$\mathfrak{Ext}^{2}\bigl{(}G,\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}\mathrel{\vcentcolon=}\operatorname{Ext}^{2}_{\mathbb{I}_{\mathrm{p}}(G)}\bigl{(}\textup{K}_{0}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}}),\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}.$
The extension group in question is described in (5.7) as the cokernel of the map
$\kappa^{*}\colon\operatorname{Hom}_{\mathbb{I}_{\mathrm{p}}(G)}\bigl{(}\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}],\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}\to\operatorname{Hom}_{\mathbb{I}_{\mathrm{p}}(G)}\bigl{(}\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}}),\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$
induced by the canonical injective diagram morphism $\kappa\colon\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}]$ in (5.5). So elements in $\mathfrak{Ext}^{2}\bigl{(}G,\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$ are represented by diagram morphisms from $\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})$ to $\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$, and such a diagram morphism represents 00 if and only if it factors through $\kappa$.
Definition 6.3.
Let $\varphi_{0}\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$ be a homomorphism. It induces a map $\psi\colon\mathbb{I}_{\mathrm{p}}(G)\cong\mathbb{I}_{\mathrm{p}}(\textup{C}^{*}(G))\to\mathbb{I}_{\mathrm{p}}(B)$. Let
$\varphi_{1}\colon\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$
be a morphism of diagrams. By Theorem 3.2, there is a correspondence $\mathcal{H}$ that lifts $\varphi_{0}$. The obstruction class $\Delta(\varphi_{0},\varphi_{1})\in\mathfrak{Ext}^{2}_{\mathbb{I}_{\mathrm{p}}(G)}\bigl{(}G,\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$ of $(\varphi_{0},\varphi_{1})$ is the image of the diagram morphism $\varphi_{1}-\mathcal{H}_{*}\colon\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$.
As the name indicates, the obstruction class obstructs the existence of a proper correspondence that lifts both maps $\varphi_{0}$ and $\varphi_{1}$:
Proposition 6.4.
The obstruction class does not depend on the choice of the proper correspondence $\mathcal{H}$ lifting $\varphi_{0}$. It vanishes if there is a proper correspondence $\mathcal{H}$ that induces the given maps $\varphi_{0}$ and $\varphi_{1}$ on $\mathcal{P}$ and $\textup{K}_{1}$.
Proof.
By definition, $\Delta(\varphi_{0},\varphi_{1})=0$ if $\mathcal{H}$ induces $\varphi_{0}$ and $\varphi_{1}$ on $\mathcal{P}$ and $\textup{K}_{1}$. So the second statement follows if the obstruction class does not depend on the choice of $\mathcal{H}$. This is what we are going to prove.
Let $\mathcal{H}^{\prime}$ be another proper correspondence lifting $\varphi_{0}$. This involves isomorphic Hilbert $B$-modules $\mathcal{H}_{v}$ because $[\mathcal{H}_{v}]=\varphi_{0}[p_{v}]$. So $\mathcal{H}^{\prime}$ is isomorphic to the proper correspondence for a family $((\mathcal{H}_{v}),(U_{v}^{\prime}))$ with the same proper Hilbert $B$-modules $\mathcal{H}_{v}$. By Lemma 3.5, the unitaries $U_{v}^{\prime}$ are of the form $U^{\prime}_{v}=\Upsilon_{v}U_{v}$ with $\Upsilon_{v}\in\mathcal{U}(\mathbb{K}(\mathcal{H}_{v}))$ for $v\in V_{\mathrm{reg}}$. Each $\Upsilon_{v}$ has a class in $\textup{K}_{1}(\mathbb{K}(\mathcal{H}_{v}))$. The $\textup{C}^{*}$-algebra $\mathbb{K}(\mathcal{H}_{v})$ is Morita–Rieffel equivalent through $\mathcal{H}_{v}$ to the ideal in $B$ generated by the inner products of elements in $\mathcal{H}_{v}$. This ideal is equal to $\psi(\langle v\rangle)\mathrel{\triangleleft}B$, where $\langle v\rangle\subseteq V$ is the hereditary and saturated subset generated by $v$. Thus $\mathcal{H}^{\prime}$ yields a diagram morphism
(6.5) $[\Upsilon]\colon\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}]\to\psi^{*}(\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})),\qquad\delta_{v}\mapsto[\Upsilon_{v}],$
compare the proof of Proposition 5.6. We claim that
(6.6) $\mathcal{H}_{*}-\mathcal{H}^{\prime}_{*}=[\Upsilon]\circ\kappa\qquad\text{as maps }\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$
with the map $\kappa$ in (5.5). So $\mathcal{H}$ and $\mathcal{H}^{\prime}$ have the same image in $\mathfrak{Ext}^{2}\bigl{(}G,\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$, as desired. To prove (6.6), choose a hereditary and saturated subset $W\subseteq V$ and $x\in\textup{K}_{1}(\textup{C}^{*}(G)_{W})$. Write $\kappa(x)=\sum_{v\in W_{\mathrm{reg}}}{}(c_{v}^{+}-c_{v}^{-})\delta_{v}$ with $c_{v}^{\pm}\geq 0$. Let $v\in W_{\mathrm{reg}}$. The Hilbert $B$-modules $\mathcal{F}_{v}\mathrel{\vcentcolon=}\bigoplus_{v\in V}\mathcal{H}_{v}^{\oplus c_{v}^{+}}\oplus\bigoplus_{e\in E^{v}}\mathcal{H}_{s(e)}^{\oplus c_{v}^{-}}$ and $\bigoplus_{v\in V}\mathcal{H}_{v}^{\oplus c_{v}^{-}}\oplus\bigoplus_{e\in E^{v}}\mathcal{H}_{s(e)}^{\oplus c_{v}^{+}}$ are isomorphic by a permutation $\Sigma$ of the summands because $(\textup{id}-M_{G}^{W})\kappa(x)=0$. Since their inner products take values in $\psi(W)\mathrel{\triangleleft}B$, they are Hilbert $\psi(W)$-modules. The unitaries
$U_{x}$ :$\mathrel{\vcentcolon=}\bigoplus_{v\in V_{\mathrm{reg}}}U_{v}^{\oplus c_{v}^{+}}\oplus\bigoplus_{v\in V_{\mathrm{reg}}}(U_{v}^{*})^{\oplus c_{v}^{-}},$
$U^{\prime}_{x}$ :$\mathrel{\vcentcolon=}\bigoplus_{v\in V_{\mathrm{reg}}}(\Upsilon_{v}U_{v})^{\oplus c_{v}^{+}}\oplus\bigoplus_{v\in V_{\mathrm{reg}}}(U_{v}^{*}\Upsilon_{v}^{*})^{\oplus c_{v}^{-}}$
are isomorphisms between these Hilbert $\psi(W)$-modules as well. So $\Sigma U_{x},\Sigma U_{x}^{\prime}\in\mathcal{U}(\mathcal{F}_{v})$. By definition, $\mathcal{H}_{*}(x)$ and $\mathcal{H}^{\prime}_{*}(x)$ are the classes in $\textup{K}_{1}(\psi(W))$ defined by the unitaries $\Sigma U_{x}$ and $\Sigma U_{x}^{\prime}$. They differ by the class of $\bigoplus_{v\in V_{\mathrm{reg}}}\Upsilon_{v}^{\oplus c_{v}^{+}}\oplus\bigoplus_{v\in V_{\mathrm{reg}}}(\Upsilon_{v}^{*})^{\oplus c_{v}^{-}}$. This is $[\Upsilon](\kappa(x))$ because $[\Upsilon_{v}^{*}]=-[\Upsilon_{v}]$ in $\textup{K}_{1}(\psi(W))$. ∎
The vanishing of the obstruction class is necessary for the existence of a proper correspondence $\mathcal{H}$ that lifts $(\varphi_{0},\varphi_{1})$. We are going to show that it is also sufficient under an extra assumption on $B$.
Definition 6.7.
A unital $\textup{C}^{*}$-algebra $B$ is $\textup{K}_{1}$-surjective or $\textup{K}_{1}$-injective, respectively, if the map $\pi_{0}(\mathcal{U}(B))\to\textup{K}_{1}(B)$ is surjective or injective, respectively.
Proposition 6.8.
Let $G$ be a graph, $B$ a $\textup{C}^{*}$-algebra, and let $\varphi_{0}\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$ and $\varphi_{1}\colon\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$ be given, where $\psi\colon\mathbb{I}_{\mathrm{p}}(G)\to\mathbb{I}_{\mathrm{p}}(B)$ is induced by $\varphi_{0}$. For $v\in V$, let $\mathcal{H}_{v}$ be proper Hilbert $B$-modules with $[\mathcal{H}_{v}]=\varphi_{0}[p_{v}]$. Assume that the unital $\textup{C}^{*}$-algebras $\mathbb{K}(\mathcal{H}_{v})$ for $v\in V$ are $\textup{K}_{1}$-surjective. There is a proper $\textup{C}^{*}(G),B$-correspondence $\mathcal{H}$ that induces the given maps $\varphi_{0}$ and $\varphi_{1}$ if and only if $\Delta(\varphi_{0},\varphi_{1})=0$.
Proof.
One direction is proven already in Proposition 6.4, without assumption on $B$. Here we prove the converse. So assume that the obstruction class vanishes. We shall modify a proper $\textup{C}^{*}(G),B$-correspondence $\mathcal{H}$ that lifts $\varphi_{0}$ so that it also lifts $\varphi_{1}$. The vanishing of the obstruction class means that
$\varphi_{1}-\mathcal{H}_{*}=\alpha\circ\kappa\colon\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$
for a diagram morphism $\alpha\colon\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}]\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$. Describe $\mathcal{H}$ by a family $((\mathcal{H}_{v}),(U_{v}))$ as above. Let $v\in V_{\mathrm{reg}}$ and let $\langle v\rangle\subseteq V$ be the hereditary and saturated subset generated by $v$. Since $\psi(\langle v\rangle)\mathrel{\triangleleft}B$ is Morita–Rieffel equivalent to $\mathbb{K}(\mathcal{H}_{v})$ and $\mathbb{K}(\mathcal{H}_{v})$ is $\textup{K}_{1}$-surjective, $\alpha(\delta_{v})\in\textup{K}_{1}(\psi(\langle v\rangle))\cong\textup{K}_{1}(\mathbb{K}(\mathcal{H}_{v}))$ is represented by some $\Upsilon_{v}\in\mathcal{U}(\mathbb{K}(\mathcal{H}_{v}))$.
Define $U_{v}^{\prime}\mathrel{\vcentcolon=}\Upsilon_{v}\circ U_{v}$. The family $((\mathcal{H}_{v}),(U_{v}^{\prime}))$ gives another proper $\textup{C}^{*}(G),B$-correspondence $\mathcal{H}^{\prime}$ lifting the same map $\varphi_{0}$. Let $[\Upsilon_{v}]$ denote the class of $\Upsilon_{v}$ in $\textup{K}_{1}(\psi(\langle v\rangle))\cong\textup{K}_{1}(\mathbb{K}(\mathcal{H}_{v}))$. As above, this defines a morphism of diagrams $\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}]\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$. The proof of Proposition 6.4 shows that $\mathcal{H}^{\prime}_{*}-\mathcal{H}_{*}$ is equal to $[\Upsilon_{v}]\circ\kappa=\alpha\circ\kappa=\varphi_{1}-\mathcal{H}_{*}$. As a result, $\mathcal{H}^{\prime}_{*}=\varphi_{1}$. ∎
Now assume that the target $B$ is a graph $\textup{C}^{*}$-algebra as well. The following theorem allows us to apply all of the previous results to study $A,B$-correspondences:
Theorem 6.9.
Let $G_{2}=(r,s\colon E_{2}\rightrightarrows V_{2})$ be a regular directed graph and let $B$ be its graph $\textup{C}^{*}$-algebra. Let $p\in\mathcal{P}(B)$. Then $B$ has stable weak cancellation and the unital $\textup{C}^{*}$-algebra $p(B\otimes\mathbb{K})p$ is $\textup{K}_{1}$-bijective.
Proof.
Graph $\textup{C}^{*}$-algebras have stable weak cancellation by [Ara-Moreno-Pardo:Nonstable_K_for_Graph_Algs]*Corollary 7.2. By Theorem 3.1, any $p\in\mathcal{P}(B)$ is equivalent to $\sum_{v\in V_{2}}c_{v}[p_{v}]$ for a finitely supported function $(c_{v})_{v\in V_{2}}\in\mathbb{N}[V]$.
For a finite subset $V^{\prime}_{2}\subseteq V_{2}$, let $B^{\prime}$ be the $\textup{C}^{*}$-subalgebra of $B$ generated by $\{t_{e}\,{:}\,\mathopen{}r(e)\in V^{\prime}_{2}\}\cup\{p_{v}\,{:}\,\mathopen{}v\in V^{\prime}_{2}\}$. These form a directed set of $\textup{C}^{*}$-subalgebras of $B$, and their union is dense in $B$. Therefore, $p(B\otimes\mathbb{K})p$ is the inductive limit of the $\textup{C}^{*}$-algebras $p(B^{\prime}\otimes\mathbb{K})p$. Both $\textup{K}_{1}$ and the set of connected components of the unitary group are continuous for inductive limits. Therefore, it is enough to prove that $p(B^{\prime}\otimes\mathbb{K})p$ is $\textup{K}_{1}$-bijective for all finite subsets $V^{\prime}_{2}\subseteq V_{2}$. We may assume without loss of generality that $V^{\prime}_{2}$ contains $\operatorname{supp}(c_{v})$, so that $p\in B^{\prime}\otimes\mathbb{K}\subseteq B\otimes\mathbb{K}$. The $\textup{C}^{*}$-algebra $B^{\prime}$ may be identified with the $\textup{C}^{*}$-algebra of a finite graph (compare [Raeburn-Szymanski:CK_algs_of_inf_graphs_and_matrices]*Lemma 1.2). Here we may take the graph with vertex set $V^{\prime}_{2}\cup s(r^{-1}(V^{\prime}_{2}))$ and edge set $r^{-1}(V^{\prime}_{2})$. Therefore, it suffices to prove that $p(B\otimes\mathbb{K})p$ is $\textup{K}_{1}$-bijective when $B$ is the graph $\textup{C}^{*}$-algebra of a finite graph. We assume this from now on.
The advantage in the finite case is that there are only finitely many gauge-invariant ideals. Therefore, there is an increasing chain of gauge-invariant ideals $0=I_{0}\subset I_{1}\subset I_{2}\subset\dotsb\subset I_{\ell}=B$ so that each subquotient $I_{k+1}/I_{k}$ for $k=0,\dotsc,\ell-1$ has no gauge-invariant ideals. It is known that these subquotients $I_{k+1}/I_{k}$ are again Morita equivalent to graph $\textup{C}^{*}$-algebras, so that they have stable weak cancellation. In addition, they are either AF or of the form $\mathrm{C}(\mathbb{T},\mathbb{M}_{n})$ for some $n\in\mathbb{N}_{\geq 1}$ or purely infinite simple. In the first two cases, $p\cdot(I_{k+1}/I_{k}\otimes\mathbb{K})p$ has stable rank $1$ (see [Raeburn-Szymanski:CK_algs_of_inf_graphs_and_matrices]*Proposition 5.4). In all three cases, $p\cdot(I_{k+1}/I_{k}\otimes\mathbb{K})\cdot p$ is extremally rich and has weak cancellation (see [Brown-Pedersen:Non-stable]). Brown and Pedersen show in [Brown-Pedersen:Non-stable] that extremally rich $\textup{C}^{*}$-algebras with weak cancellation have the following properties:
  • $\textup{K}_{1}$-surjectivity by [Brown-Pedersen:Non-stable]*Theorem 4.4;
  • good index theory by [Brown-Pedersen:Non-stable]*Theorem 5.1;
  • $\textup{K}_{1}$-injectivity and weak $\textup{K}_{0}$-surjectivity by [Brown-Pedersen:Non-stable]*Theorem 6.7.
It is shown in [Brown-Pedersen:Non-stable] that the class of $\textup{C}^{*}$-algebras with these four properties – weak $\textup{K}_{0}$-surjectivity, $\textup{K}_{1}$-surjectivity, $\textup{K}_{1}$-injectivity, and good index theory – is closed under extensions (see [Brown-Pedersen:Non-stable]*Propositions 6.6 and 7.2.4). Therefore, $p(B\otimes\mathbb{K})p$ has these four properties. In particular, it is $\textup{K}_{1}$-bijective. ∎
7. The purely infinite case
At this point, our results combined with Kirchberg’s Classification Theorem imply a criterion when there is a $\textup{C}^{*}$-algebra isomorphism $\textup{C}^{*}(G)\cong B$, even without assuming $B$ to be a graph $\textup{C}^{*}$-algebra:
Theorem 7.1.
Let $G$ be a countable graph for which $\textup{C}^{*}(G)$ is purely infinite. Let $B$ be a separable, nuclear, strongly purely infinite $\textup{C}^{*}$-algebra such that all ideals of $B$ are in the bootstrap class and such that $\operatorname{Prim}(B)$ has a basis consisting of quasi-compact open subsets. Assume that there are an isomorphism $\psi\colon\mathbb{I}_{\mathrm{p}}(G)\to\mathbb{I}_{\mathrm{p}}(B)$ of partially ordered sets and isomorphisms of diagrams $\varphi_{j}\colon\textup{K}_{j}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\xrightarrow{\sim}\psi^{*}\textup{K}_{j}(B,\mathbb{I}_{\mathrm{p}})$ for $j=0,1$ for which the obstruction class vanishes. Then $\textup{C}^{*}(G)$ and $B$ are stably isomorphic, and vice versa.
The proof of the converse direction (“vice versa”) is easy. The proof of the interesting direction requires some preparatory results.
Lemma 7.2.
Let $X$ be a topological space. Let $\mathbb{CO}(X)\subseteq\mathbb{O}(X)$ be the set of compact-open subsets. Assume that $\mathbb{CO}(X)$ is a basis. Let $\mathcal{F}(X)$ denote the set of subsets of $\mathbb{CO}(X)$ which are closed under finite unions and subsets. Then the map $U\mapsto F_{U}\mathrel{\vcentcolon=}\{V\in\mathbb{CO}(X)\mid V\subseteq U\}$ is an isomorphism of partially ordered sets $\mathbb{O}(X)\cong\mathcal{F}(X)$ with inverse $S\mapsto\bigcup S$.
Proof.
The maps $F\colon\mathbb{O}(X)\to\mathcal{F}(X)$ and ${\bigcup}\colon\mathcal{F}(X)\to\mathbb{O}(X)$ are well defined monotone maps. One composite is the identity as $U=\bigcup{}\{V\in\mathbb{CO}(X)\,{:}\,\mathopen{}V\subseteq U\}$ for all $U\in\mathbb{O}(X)$ by assumption on $X$; the other is the identity because
$S\subseteq\{V\in\mathbb{CO}(X)\,{:}\,\mathopen{}V\subseteq\bigcup S\}\subseteq S$
for all $S\in\mathcal{F}(X)$, where the first inclusion is obvious and the second is seen as follows: if $V\subseteq\bigcup S$ is compact, then $V\subseteq\bigcup T$ for some finite subset $T\subseteq S$ and then $V\in S$ by assumption on $S$. ∎
Proposition 7.3.
Let $B$ be a separable purely infinite $\textup{C}^{*}$-algebra such that the topology on $\operatorname{Prim}(B)$ has a basis consisting of compact-open subsets. Let $\mathcal{F}(\mathbb{I}_{\mathrm{p}}(B))$ denote the set of subsets of $\mathbb{I}_{\mathrm{p}}(B)$ which are closed under finite unions and subsets. Then $I\mapsto F_{I}\mathrel{\vcentcolon=}\{J\in\mathbb{I}_{\mathrm{p}}(B)\mid J\subseteq I\}$ defines an isomorphism of partially ordered sets $\mathbb{I}(B)\cong\mathcal{F}(\mathbb{I}_{\mathrm{p}}(B))$ with inverse $S\mapsto\overline{\sum S}$.
Proof.
[Pasnicu-Rordam:Purely_infinite_rr0]*Proposition 2.7 says that $\mathbb{I}_{\mathrm{p}}(B)$ is the set of all ideals $I\mathrel{\triangleleft}B$ for which $\widehat{I}$ is quasi-compact. Now we can deduce the result from Lemma 7.2, using the isomorphism $\mathbb{O}\bigl{(}\operatorname{Prim}(A)\bigr{)}\cong\mathbb{I}(A)$, $U\mapsto\bigcap\operatorname{Prim}(A)\setminus U$ (see [Dixmier:Cstar-algebres]*§3.2). ∎
Proof of Theorem 7.1.
Since $\textup{C}^{*}(G)$ and $B$ are purely infinite, Corollary 4.16 shows that an isomorphism $\varphi_{0}\colon\textup{K}_{0}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\xrightarrow{\sim}\psi^{*}\textup{K}_{0}(B,\mathbb{I}_{\mathrm{p}})$ is equivalent to an isomorphism $\mathcal{P}(\textup{C}^{*}(G))\cong\mathcal{P}(B)$. Then the obstruction class in the statement is well defined. If it vanishes, Proposition 6.8 gives a proper $\textup{C}^{*}(G),B$-correspondence $\mathcal{H}$ that induces the given isomorphisms $\varphi_{0}$ and $\varphi_{1}$. The proper correspondence $\mathcal{H}$ has a class in the equivariant KK-group $\textup{KK}(\operatorname{Prim}\textup{C}^{*}(G);\textup{C}^{*}(G),B)$. It induces an isomorphism on the K-theory of all ideals in $\mathbb{I}_{\mathrm{p}}(B)\cong\mathbb{I}_{\mathrm{p}}(\textup{C}^{*}(G))$. By the Universal Coefficient Theorem, it induces E-theory equivalences on these ideals.
Every ideal in $\textup{C}^{*}(G)$ or in $B$ is generated by its projections. This follows for $\textup{C}^{*}(G)$ because every ideal in a purely infinite graph $\textup{C}^{*}$-algebra is gauge-invariant and for $B$ by [Pasnicu-Rordam:Purely_infinite_rr0]*Proposition 2.7. It also follows from [Pasnicu-Rordam:Purely_infinite_rr0]*Proposition 2.7 that $\operatorname{Prim}(\textup{C}^{*}(G))$ has a basis consisting of quasi-compact open subsets. Hence Proposition 7.3 shows that $\psi$ extends canonically to a lattice isomorphism $\mathbb{I}(\textup{C}^{*}(G))\xrightarrow{\sim}\mathbb{I}(B)$. Thus $\operatorname{Prim}(\textup{C}^{*}(G))\cong\operatorname{Prim}(B)$. The quasi-compact open subsets form a basis for the topology on these primitive ideal spaces, and the correspondence $\mathcal{H}$ induces E-theory equivalences on those ideals that correspond to quasi-compact open subsets. Thus [Dadarlat-Meyer:E_over_space]*Theorems 3.2 and 3.10 imply that $\mathcal{H}$ induces a $\operatorname{Prim}\textup{C}^{*}(G)$-equivariant E-theory equivalence. Then the corresponding equivariant KK-class is also invertible by [Gabe:Lifting_theorems_for_cp_maps]*Theorem 6.2. Hence Kirchberg’s Classification Theorem allows to lift it to a stable isomorphism between $\textup{C}^{*}(G)$ and $B$. ∎
Remark 7.4.
The constructed stable isomorphism $\textup{C}^{*}(G)\otimes\mathbb{K}\cong B\otimes\mathbb{K}$ also gives a proper $\textup{C}^{*}(G),B$-correspondence. So it corresponds to a particular choice of the proper Hilbert $B$-modules $\mathcal{H}_{v}$ and the unitaries $U_{v}$ in our description of such correspondences. Of course, it is very hard to find these unitaries directly.
If the target $\textup{C}^{*}$-algebra $B$ is known to be a graph $\textup{C}^{*}$-algebra as well, then the proof of Theorem 7.1 simplifies considerably. Then Theorem 8.7 and Proposition 7.3 show that the $\textup{C}^{*}$-correspondence above is an ideal-preserving stable homotopy equivalence. This forces it to be an ideal-preserving KK-equivalence.
8. Homotopy classes of correspondences
In this section, we first study when two proper $\textup{C}^{*}(G),B$-correspondences $\mathcal{H}$ and $\mathcal{H}^{\prime}$ are homotopic. Secondly, we use this to prove a criterion for two graph $\textup{C}^{*}$-algebras to be homotopy equivalent.
Lemma 8.1.
Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be homotopic. Then they induce the same maps
$\varphi_{0}\colon\mathcal{P}(\textup{C}^{*}(G))$ $\to\mathcal{P}(B),$ $\qquad\psi\colon\mathbb{I}_{\mathrm{p}}^{\infty}(\textup{C}^{*}(G))$ $\to\mathbb{I}_{\mathrm{p}}^{\infty}(B),$
$\varphi_{1}\colon\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})$ $\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}}).$
The entire homotopy $(\mathcal{H}_{t})_{t\in[0,1]}$ preserves all ideals in $\mathbb{I}_{\mathrm{p}}^{\infty}(\textup{C}^{*}(G))$ in the sense that each $\mathcal{H}_{t}$ induces the same map $\mathbb{I}_{\mathrm{p}}^{\infty}(\textup{C}^{*}(G))\to\mathbb{I}_{\mathrm{p}}^{\infty}(B)$.
Proof.
The maps $\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B)$ induced by $\mathcal{H}$ and $\mathcal{H}^{\prime}$ are the same because homotopic projections are Murray–von Neumann equivalent. Therefore, $\mathcal{H}$ and $\mathcal{H}^{\prime}$ – and also $\mathcal{H}_{t}$ for $t\in[0,1]$ – induce the same map $\mathbb{I}_{\mathrm{p}}^{\infty}(\textup{C}^{*}(G))\to\mathbb{I}_{\mathrm{p}}^{\infty}(B)$. Since $\textup{K}_{1}$ is defined using homotopy of unitaries, it follows that the diagram morphisms $\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$ induced by $\mathcal{H}$ and $\mathcal{H}^{\prime}$ are equal as well. Even more, they are constant along the homotopy $(\mathcal{H}_{t})$. ∎
Let $\mathcal{H}$ and $\mathcal{H}^{\prime}$ be $\textup{C}^{*}(G),B$-correspondences that induce the same maps
$\varphi_{0}\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B),\qquad\varphi_{1}\colon\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}}).$
The first condition implies that $\mathcal{H}$ and $\mathcal{H}^{\prime}$ also induce the same map $\psi\colon\mathbb{I}_{\mathrm{p}}(\textup{C}^{*}(G))\to\mathbb{I}_{\mathrm{p}}(B)$, which is needed to formulate the second condition. Describe $\mathcal{H}$ and $\mathcal{H}^{\prime}$ by families $\bigl{(}(\mathcal{H}_{v}),(U_{v})\bigr{)}$ and $\bigl{(}(\mathcal{H}_{v}^{\prime}),(U^{\prime}_{v})\bigr{)}$, respectively, as in Theorem 3.2. Since $[\mathcal{H}_{v}]=[\mathcal{H}_{v}^{\prime}]$ in $\mathcal{P}(B)$, the Hilbert $B$-modules $\mathcal{H}_{v}$ and $\mathcal{H}^{\prime}_{v}$ are isomorphic. For simplicity, we replace $\mathcal{H}^{\prime}$ by an isomorphic correspondence so that these Hilbert modules become equal. Then Lemma 3.5 gives unitaries $\Upsilon_{v}\in\mathbb{K}(\mathcal{H}_{v})$ for $v\in V_{\mathrm{reg}}$ with $U_{v}^{\prime}=\Upsilon_{v}U_{v}$.
Next, we describe an obstruction for $\mathcal{H}$ and $\mathcal{H}^{\prime}$ to be homotopic. This obstruction will turn out to be sharp under an extra assumption on $B$. Let $v\in V_{\mathrm{reg}}$. Let $\langle v\rangle\in\mathbb{I}_{\mathrm{p}}(G)$ be the hereditary and saturated subset generated by $v$. Since this corresponds to the ideal in $\textup{C}^{*}(G)$ generated by $p_{v}$, it follows that $\psi(\langle v\rangle)\mathrel{\triangleleft}B$ is the ideal generated by the matrix coefficients of vectors in $\mathcal{H}_{v}$. Thus $\mathcal{H}_{v}$ is a Morita–Rieffel equivalence between $\mathbb{K}(\mathcal{H}_{v})$ and $\psi(\langle v\rangle)$. As in the proof of Proposition 6.8, we let $[\Upsilon_{v}]$ denote the class of $\Upsilon_{v}$ in $\textup{K}_{1}(\psi(\langle v\rangle))\cong\textup{K}_{1}(\mathbb{K}(\mathcal{H}_{v}))$. We combine these classes for $v\in V_{\mathrm{reg}}$ into a morphism of diagrams $[\Upsilon]\colon\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}]\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$ as in (6.5). Since $\mathcal{H}$ and $\mathcal{H}^{\prime}$ induce the same map $\varphi_{1}$, (6.6) implies $[\Upsilon]\circ\kappa=0$.
We have described the group
$\mathfrak{Ext}^{1}\bigl{(}G,\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}\mathrel{\vcentcolon=}\operatorname{Ext}^{1}_{\mathbb{I}_{\mathrm{p}}(G)}\bigl{(}\textup{K}_{0}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}}),\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$
as the first cohomology group of the cochain complex in (5.7). The equation $[\Upsilon]\circ\kappa=0$ says that $[\Upsilon]$ is a cocycle and thus defines a class in $\mathfrak{Ext}^{1}\bigl{(}G,\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$.
Proposition 8.2.
If $\mathcal{H}$ and $\mathcal{H}^{*}$ are homotopic $\textup{C}^{*}(G),B$-correspondences, then $[\Upsilon]$ becomes 00 in $\mathfrak{Ext}^{1}\bigl{(}G,\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$. In particular, $[\Upsilon]$ does not depend on the choices made in its construction.
Proof.
By the remarks after Lemma 3.5, $\mathcal{H}$ and $\mathcal{H}^{\prime}$ are homotopic if and only if we may find unitaries $W_{v,0},W_{v,1}\in\mathcal{U}(\mathcal{H}_{v})$ for all $v\in V_{\mathrm{reg}}$ so that
$\Upsilon_{v,0}$ :$\mathrel{\vcentcolon=}W_{v,0}\circ U_{v}\circ\biggl{(}\bigoplus_{e\in E^{v}}W_{s(e),0}^{*}\biggr{)}\circ U_{v}^{*},$
$\Upsilon_{v,1}$ :$\mathrel{\vcentcolon=}W_{v,1}\circ\Upsilon_{v}U_{v}\circ\biggl{(}\bigoplus_{e\in E^{v}}W_{s(e),1}^{*}\biggr{)}\circ U_{v}^{*}$
are homotopic in $\mathcal{U}(\mathcal{H}_{v})$. Equivalently, $\Upsilon_{v}$ is homotopic to
(8.3) $W_{v}\cdot U_{v}\biggl{(}\bigoplus_{e\in E^{v}}W_{s(e)}^{*}\biggr{)}U_{v}^{*},$
where $W_{v}\mathrel{\vcentcolon=}W_{v,1}^{*}W_{v,0}$. This is true by assumption. In the same way that the unitaries $\Upsilon_{v}\in\mathcal{U}(\mathcal{H}_{v})$ for $v\in V_{\mathrm{reg}}$ give a map $[\Upsilon]\colon\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}]\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$, the unitaries $W_{v}$ give a map $[W]\colon\mathbb{Z}[V,\mathbb{I}_{\mathrm{p}}]\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$. The class of the unitary in (8.3) in $\textup{K}_{1}(\mathbb{K}(\mathcal{H}_{v}))\cong\textup{K}_{1}(\psi(W_{v}))$ is $[W_{v}]-\sum_{e\in E^{v}}{}[W_{s(e)}]$. Therefore, $[W]\circ(\textup{id}-M_{G}^{\mathbb{I}_{\mathrm{p}}})=[\Upsilon]$ as diagram morphisms $\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}]\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$. This witnesses that $[\Upsilon]$ represents zero in $\mathfrak{Ext}^{1}\bigl{(}G,\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$. ∎
Theorem 8.4.
Describe a $\textup{C}^{*}(G),B$-correspondences $\mathcal{H}$ by a family $\bigl{(}(\mathcal{H}_{v}),(U_{v})\bigr{)}$ as in Theorem 3.2. Let
$\varphi_{0}\colon\mathcal{P}(\textup{C}^{*}(G))\to\mathcal{P}(B),\qquad\varphi_{1}\colon\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$
be the maps induced by $\mathcal{H}$. Assume that the unital $\textup{C}^{*}$-algebras $\mathbb{K}(\mathcal{H}_{v})$ for $v\in V_{\mathrm{reg}}$ are $\textup{K}_{1}$-bijective. Then the construction in Proposition 8.2 gives a bijection between the group $\mathfrak{Ext}^{1}\bigl{(}G,\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$ and the set of homotopy classes of $\textup{C}^{*}(G),B$-correspondences that induce the same maps $\varphi_{0}$ and $\varphi_{1}$.
Proof.
Let $\mathcal{H}^{\prime}$ and $\mathcal{H}^{\prime\prime}$ be $\textup{C}^{*}(G),B$-correspondences that induce the same maps $\varphi_{0}$ and $\varphi_{1}$ as $\mathcal{H}$. Describe $\mathcal{H}^{\prime}$ and $\mathcal{H}^{\prime\prime}$ by families $\bigl{(}(\mathcal{H}_{v}^{\prime}),(U^{\prime}_{v})\bigr{)}$ and $\bigl{(}(\mathcal{H}_{v}^{\prime\prime}),(U^{\prime\prime}_{v})\bigr{)}$ and replace them by isomorphic correspondences with $\mathcal{H}^{\prime\prime}_{v}=\mathcal{H}^{\prime}_{v}=\mathcal{H}_{v}$ for all $v\in V$. Then there are unitaries $\Upsilon_{v}^{\prime}$ and $\Upsilon_{v}^{\prime\prime}$ for $v\in V_{\mathrm{reg}}$ with $U_{v}^{\prime}=\Upsilon^{\prime}_{v}U_{v}$ and $U_{v}^{\prime\prime}=\Upsilon^{\prime\prime}_{v}U_{v}$. Assume that $[\Upsilon^{\prime}]=[\Upsilon^{\prime\prime}]$ holds in $\mathfrak{Ext}^{1}\bigl{(}G,\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$. Then there is a map $\beta\colon\mathbb{Z}[V,\mathbb{I}_{\mathrm{p}}]\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$ with $\beta\circ(\textup{id}-M_{G}^{\mathbb{I}_{\mathrm{p}}})=[\Upsilon^{\prime\prime}]-[\Upsilon^{\prime}]=[(\Upsilon^{\prime})^{*}\Upsilon^{\prime\prime}]$. The proof of Proposition 6.8 shows that $\beta$ is specified by elements $\beta_{v}\in\textup{K}_{1}(\psi(\langle v\rangle))\cong\textup{K}_{1}(\mathbb{K}(\mathcal{H}_{v}))$ for $v\in V$. Since $\mathbb{K}(\mathcal{H}_{v})$ is $\textup{K}_{1}$-surjective, these $\textup{K}_{1}$-classes are already represented by unitaries $W_{v}\in\mathcal{U}(\mathcal{H}_{v})$. The equation $\beta\circ(\textup{id}-M_{G}^{\mathbb{I}_{\mathrm{p}}})=[(\Upsilon^{\prime})^{*}\Upsilon^{\prime\prime}]$ says that the unitaries in (8.3) and $(\Upsilon_{v}^{\prime})^{*}\Upsilon_{v}^{\prime\prime}$ represent the same class in $\textup{K}_{1}(\psi(\langle v\rangle))\cong\textup{K}_{1}(\mathbb{K}(\mathcal{H}_{v}))$. Since $\mathbb{K}(\mathcal{H}_{v})$ is $\textup{K}_{1}$-injective, $(\Upsilon_{v}^{\prime})^{*}\Upsilon_{v}^{\prime\prime}$ and $W_{v}\cdot U_{v}\biggl{(}\bigoplus_{e\in E^{v}}W_{s(e)}^{*}\biggr{)}U_{v}^{*}$ are homotopic in $\mathbb{K}(\mathcal{H}_{v})$. Similarly, $\Upsilon^{\prime}_{v}W_{v}$ is homotopic to $W_{v}\Upsilon^{\prime}_{v}$. So $\Upsilon_{v}^{\prime\prime}$ is homotopic in $\mathbb{K}(\mathcal{H}_{v})$ to $W_{v}\cdot\Upsilon^{\prime}_{v}U_{v}\biggl{(}\bigoplus_{e\in E^{v}}W_{s(e)}^{*}\biggr{)}U_{v}^{*}$. The latter family together with $\mathcal{H}_{v}$ for $v\in V$ describes a $\textup{C}^{*}(G),B$-correspondence that is isomorphic to $\mathcal{H}^{\prime}$, and the homotopy of unitaries gives a homotopy between the latter and $\mathcal{H}^{\prime\prime}$. So $\mathcal{H}^{\prime}$ and $\mathcal{H}^{\prime\prime}$ are homotopic.
Next, take any class in $\mathfrak{Ext}^{1}\bigl{(}G,\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$. This is represented by some diagram morphism $\upsilon\colon\mathbb{Z}[V_{\mathrm{reg}},\mathbb{I}_{\mathrm{p}}]\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$ with $\upsilon\circ\kappa=0$. For $v\in V_{\mathrm{reg}}$, the value of this map on $\delta_{v}\in\mathbb{Z}[\langle v\rangle]$ is a class in $\textup{K}_{1}(\psi(\langle v\rangle))\cong\textup{K}_{1}(\mathbb{K}(\mathcal{H}_{v}))$. The latter is represented by a unitary $\Upsilon_{v}\in\mathbb{K}(\mathcal{H}_{v})$ because $\mathbb{K}(\mathcal{H}_{v})$ is $\textup{K}_{1}$-bijective. The family $\bigl{(}(\mathcal{H}_{v}),(\Upsilon_{v}U_{v})\bigr{)}$ describes a $\textup{C}^{*}(G),B$-correspondence $\mathcal{H}^{\prime}$ that induces the same map $\varphi_{0}$. Since $\upsilon\circ\kappa=0$, it also induces the same diagram morphism $\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})\to\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$. By construction, $\mathcal{H}^{\prime}$ reproduces the class in $\mathfrak{Ext}^{1}\bigl{(}G,\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$ that we started with. ∎
Now let $G$ and $G^{\prime}$ be two graphs. We denote the sets of vertices and edges of $G,G^{\prime}$ by $V,V^{\prime}$ and $E,E^{\prime}$, respectively. Let $A\mathrel{\vcentcolon=}\textup{C}^{*}(G)$ and $B\mathrel{\vcentcolon=}\textup{C}^{*}(G^{\prime})$. Assume that there are an order isomorphism $\psi\colon\mathbb{I}_{\mathrm{p}}(G)\cong\mathbb{I}_{\mathrm{p}}(A)\xrightarrow{\sim}\mathbb{I}_{\mathrm{p}}(B)\cong\mathbb{I}_{\mathrm{p}}(G^{\prime})$ and diagram isomorphisms
$\varphi_{0}^{+}\colon\textup{K}_{0}^{+}(A,\mathbb{I}_{\mathrm{p}})$ $\xrightarrow{\sim}\psi^{*}\textup{K}_{0}^{+}(B,\mathbb{I}_{\mathrm{p}}),$
$\varphi_{1}\colon\textup{K}_{1}(A,\mathbb{I}_{\mathrm{p}})$ $\xrightarrow{\sim}\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}}).$
Since graph $\textup{C}^{*}$-algebras have stable weak cancellation by Theorem 6.9 and monoid isomorphisms preserve order units, $\psi$ and $\varphi_{0}^{+}$ come from a monoid isomorphism
$\varphi_{0}\colon\mathcal{P}(A)\xrightarrow{\sim}\mathcal{P}(B)$
by Theorem 4.12. When is there a stable homotopy equivalence between $A$ and $B$ that induces the pair of maps $\varphi_{0}$ and $\varphi_{1}$? In order to lift this pair of maps to proper correspondences, we need the obstruction classes of $(\varphi_{0},\varphi_{1})$ and $(\varphi_{0}^{-1},\varphi_{1}^{-1})$ to vanish. We are going to prove that this condition is also sufficient for a homotopy equivalence. First, we prove that one of these two obstruction classes vanishes if and only if the other does so.
Lemma 8.5.
The obstruction class of $(\varphi_{0},\varphi_{1})$ vanishes if and only if the obstruction class of $(\varphi_{0}^{-1},\varphi_{1}^{-1})$ vanishes.
Proof.
By symmetry, it suffices to prove one implication. So assume that the obstruction class of $(\varphi_{0},\varphi_{1})$ vanishes. Then Proposition 6.8 gives an $A,B$-correspondence $\mathcal{H}_{AB}$ that lifts $(\varphi_{0},\varphi_{1})$ because graph $\textup{C}^{*}$-algebras are $\textup{K}_{1}$-surjective by Theorem 6.9. Theorem 3.2 gives a $B,A$-correspondence $\mathcal{H}_{BA}$ that lifts $\varphi_{0}^{-1}$, but we do not yet control the map on $\textup{K}_{1}$ induced by $\mathcal{H}_{BA}$. The composite $B,B$-correspondence $\mathcal{H}_{AB}\otimes_{A}\mathcal{H}_{BA}$ induces the map $\varphi_{0}\circ\varphi_{0}^{-1}=\textup{id}_{\mathcal{P}(B)}$ on $\mathcal{P}(B)$. So we may use it to compute the obstruction class for the maps $\textup{id}_{\mathcal{P}(B)}$ and $\textup{id}_{\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})}$. We may also use the identity correspondence on $B$ to compute this obstruction class. Since the obstruction class is well defined by Proposition 6.4, both computations give the same result. The first one gives $\textup{id}_{\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})}-\varphi_{1}\circ\mathcal{H}_{BA,*}$ because $\mathcal{H}_{AB,*}=\varphi_{1}$ on $\textup{K}_{1}$. The second one gives just 00. Thus $\textup{id}_{\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})}-\varphi_{1}\circ\mathcal{H}_{BA,*}$ represents 00 in $\mathfrak{Ext}^{2}\bigl{(}G^{\prime},\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\bigr{)}$. This remains so when we compose with the isomorphism $\varphi_{1}^{-1}$. That is, $\varphi_{1}^{-1}-\mathcal{H}_{BA,*}\colon\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})\to(\psi^{-1})^{*}\textup{K}_{1}(A,\mathbb{I}_{\mathrm{p}})$ represents 00 in $\mathfrak{Ext}^{2}\bigl{(}G^{\prime},(\psi^{-1})^{*}\textup{K}_{1}(A,\mathbb{I}_{\mathrm{p}})\bigr{)}$. At the same time, $\varphi_{1}^{-1}-\mathcal{H}_{BA,*}$ is exactly the obstruction class of $(\varphi_{0}^{-1},\varphi_{1}^{-1})$. ∎
Lemma 8.6.
A self-correspondence $\mathcal{H}_{A}$ of a graph $\textup{C}^{*}$-algebra $A$ that induces the identity maps on $\mathcal{P}(A)$ and $\textup{K}_{1}(A,\mathbb{I}_{\mathrm{p}})$ is a stable homotopy equivalence. In addition, all maps in the homotopies involved here induce the identity map on the lattices of gauge-invariant ideals.
Proof.
Since $\mathcal{H}_{A}$ induces the identity map on $\mathcal{P}(A)$, we may assume without loss of generality that the underlying Hilbert modules $\mathcal{H}_{v}$ are $p_{v}A$ for all vertices $v$. Then $\mathcal{H}_{A}$ is described through unitaries $\Upsilon_{v}\in p_{v}Ap_{v}=\mathbb{K}(\mathcal{H}_{v})$ for $v\in V_{\mathrm{reg}}$. Namely, its representation in Theorem 3.2 involves the family of Hilbert modules $(p_{v}A)_{v\in V}$ and unitaries $(\Upsilon_{v}u_{v})_{v\in V_{\mathrm{reg}}}$ with the canonical unitaries $u_{v}$. Since $\mathcal{H}_{v}=\mathcal{H}_{v}^{\prime}=p_{v}A$, the underlying right Hilbert module of $\mathcal{H}_{A}$ is isomorphic to $\bigoplus_{v}p_{v}A=A$. Thus $\mathcal{H}_{A}$ is isomorphic to the correspondence associated to the -homomorphism $\sigma\colon A\to A$ defined by $\sigma(p_{v})=p_{v}$ for $v\in V$ and $\sigma(t_{e})=\Upsilon_{r(e)}t_{e}$ for $e\in E$.
Now let $\sigma^{\prime}\colon A\to A$ be defined by the conditions $\sigma^{\prime}(p_{v})=p_{v}$ for all $v\in V$ and $\sigma^{\prime}(t_{e})=\Upsilon_{r(e)}^{*}t_{e}$ for all $e\in E$. Then $\sigma^{\prime}\circ\sigma(p_{v})=p_{v}$ and $\sigma^{\prime}\circ\sigma(t_{e})=\Upsilon_{r(e)}^{*}\sigma(\Upsilon_{r(e)})t_{e}$. Since $\sigma$ induces the identity map on $\textup{K}_{1}(A,\mathbb{I}_{\mathrm{p}})$, the unitaries $\Upsilon_{r(e)}^{*}\sigma(\Upsilon_{r(e)})\in p_{v}Ap_{v}$ represent zero in $\textup{K}_{1}(p_{v}Ap_{v})$. Since $p_{v}\textup{C}^{*}(A)p_{v}$ is $\textup{K}_{1}$-injective for any vertex $v$ of the graph by Theorem 6.9, the unitary $\Upsilon_{r(e)}^{*}\sigma(\Upsilon_{r(e)})$ in $p_{v}Ap_{v}$ is homotopic to the unit element $p_{v}$. This gives us a homotopy between $\sigma^{\prime}\circ\sigma$ and the identity map (compare Theorem 8.4). Since $\sigma$ and $\sigma^{\prime}\circ\sigma$ induce the identity map on the invariants, it follows that $\sigma^{\prime}$ does so as well. Then we may switch the roles of $\sigma$ and $\sigma^{\prime}$ to show that $\sigma\circ\sigma^{\prime}$ is homotopic to the identity map as well. Then $\sigma$ is a homotopy equivalence. Since all gauge-invariant ideals are generated by the projections that they contain, and since all maps in the homotopies induce the same maps on $\mathcal{P}(A)$, it follows that all maps in the homotopies induce the identity map on the lattices of gauge-invariant ideals. ∎
Theorem 8.7.
Let $G$ and $G^{\prime}$ be countable graphs. Define $A\mathrel{\vcentcolon=}\textup{C}^{*}(G)$, $B\mathrel{\vcentcolon=}\textup{C}^{*}(G^{\prime})$. Let $\psi\colon\mathbb{I}_{\mathrm{p}}(G)\cong\mathbb{I}_{\mathrm{p}}(A)\xrightarrow{\sim}\mathbb{I}_{\mathrm{p}}(B)\cong\mathbb{I}_{\mathrm{p}}(G^{\prime})$ be an order isomorphism and let
$\varphi_{0}^{+}\colon\textup{K}_{0}^{+}(A,\mathbb{I}_{\mathrm{p}})$ $\xrightarrow{\sim}\psi^{*}\textup{K}_{0}^{+}(B,\mathbb{I}_{\mathrm{p}}),$
$\varphi_{1}\colon\textup{K}_{1}(A,\mathbb{I}_{\mathrm{p}})$ $\xrightarrow{\sim}\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$
be isomorphism of diagrams. These isomorphisms induce a monoid isomorphism $\varphi_{0}\colon\mathcal{P}(A)\xrightarrow{\sim}\mathcal{P}(B)$ as above. Assume that the obstruction class of $(\varphi_{0},\varphi_{1})$ vanishes. Then there is a stable homotopy equivalence between $\textup{C}^{*}(G)$ and $\textup{C}^{*}(G^{\prime})$ that induces the maps $\varphi_{0}^{+}$ and $\varphi_{1}$. The homotopies in this equivalence automatically preserve gauge-invariant ideals, that is, the induced maps on $\mathbb{I}^{\mathbb{T}}(\textup{C}^{*}(G))$ and $\mathbb{I}^{\mathbb{T}}(\textup{C}^{*}(G^{\prime}))$ are constant along the homotopies.
Proof.
Since the obstruction class of $(\varphi_{0},\varphi_{1})$ vanishes, the same holds for the inverse maps $(\varphi_{0}^{-1},\varphi_{1}^{-1})$ by Lemma 8.5. Since graph $\textup{C}^{*}$-algebras are $\textup{K}_{1}$-surjective by Theorem 6.9, Proposition 6.8 gives an $A,B$-correspondence $\mathcal{H}_{AB}$ and a $B,A$-correspondence $\mathcal{H}_{BA}$ that induce the maps $(\varphi_{0},\varphi_{1})$ and $(\varphi_{0}^{-1},\varphi_{1}^{-1})$ on the invariants. Let $\mathcal{H}_{A}\mathrel{\vcentcolon=}\mathcal{H}_{AB}\otimes_{B}\mathcal{H}_{BA}$ and $\mathcal{H}_{B}\mathrel{\vcentcolon=}\mathcal{H}_{BA}\otimes_{A}\mathcal{H}_{AB}$ be the composite $A,A$- and $B,B$-correspondences. Both induce the identity maps on $\mathcal{P}$ and on the $\textup{K}_{1}$-diagram. Thus they are stable homotopy equivalences by Lemma 8.6. Thus $\mathcal{H}_{AB}$ and $\mathcal{H}_{BA}$ are both left and right invertible up to homotopy. This forces them to be stable homotopy equivalences. In addition, the lemma shows that the homotopies implicit in this all induce the identity map on the lattice of gauge-invariant ideals. ∎
9. The circle components
The homotopy equivalence in Theorem 8.7 has the extra property that it preserves the gauge-invariant ideals. A graph $\textup{C}^{*}$-algebra may have gauge-invariant subquotients that are Morita equivalent to $\mathrm{C}(\mathbb{T})$. In this case, there are many ideals that are not gauge-invariant. In this section, we show that we may arrange the homotopy equivalence to also preserve these ideals. As before, we let $G$ be a row-finite graph. We are going to use some results of Hong and Szymański [Hong-Szymanski:Primitive_graph]. Therefore, we assume $G$ to be countable from now on. This ensures that ideals in $\textup{C}^{*}(G)$ are prime if and only if they are primitive, and this may go wrong in general (see [Abrams-Tomforde:prime_primitive, Abrams-Tomforde:Prime_primitive_graph]). The following notation is introduced in [Hong-Szymanski:Primitive_graph] to describe the primitive ideal space of $\textup{C}^{*}(G)$. Beware that [Hong-Szymanski:Primitive_graph] uses the opposite convention regarding the source and range maps on the graph.
Definition 9.1.
Write $w\leq v$ for $w,v\in V$ if there is a path in $G$ from $w$ to $v$, and $w\sim v$ if $w\leq v$ and $v\leq w$.
The relation $\leq$ is a preorder. So $\sim$ is an equivalence relation.
Definition 9.2.
A subset $M\subseteq V$ is called a maximal tail if
  1. (MT1)
    if $e\in E$ and $s(e)\in M$, then $r(e)\in M$;
  2. (MT2)
    if $v\in M\cap V_{\mathrm{reg}}$, then there is $e\in E$ with $r(e)=v$ and $s(e)\in M$;
  3. (MT3)
    for all $v,w\in M$, there is $y\in M$ with $v\geq y$ and $w\geq y$.
If $M$ is a maximal tail, let $\Omega(M)=V\setminus M$. Let $\mathcal{M}(G)$ be the set of all maximal tails in $G$. Let $\mathcal{M}_{\gamma}(G)\subseteq\mathcal{M}(G)$ be the subset of all $M\in\mathcal{M}(G)$ with the extra property that each cycle in $G|_{M}$ has an entry in $G|_{M}$. Let $\mathcal{M}_{\tau}(G)\mathrel{\vcentcolon=}\mathcal{M}(G)\setminus\mathcal{M}_{\gamma}(G)$.
Condition (MT1) holds if and only if $v\geq w$ and $w\in M$ implies $v\in M$. Conditions (MT1) and (MT2) say exactly that $\Omega(M)$ is hereditary and saturated. It is shown in [Hong-Szymanski:Primitive_graph] that the gauge-invariant ideal corresponding to a hereditary saturated subset $W\in\mathbb{I}(G)$ is prime in the complete lattice $\mathbb{I}^{\mathbb{T}}(\textup{C}^{*}(G))$ of gauge-invariant ideals if and only if $V\setminus W$ satisfies (MT3) as well as (MT1) and (MT2). Equivalently, the map $M\mapsto\Omega(M)$ is a bijection between $\mathcal{M}(G)$ and the set of prime ideals in the lattice $\mathbb{I}(G)$.
Proposition 9.3 ([Hong-Szymanski:Primitive_graph]).
Let $M\in\mathcal{M}_{\tau}(G)$. Then $M$ contains a unique simple cycle $e_{1}e_{2}\dotsm e_{k}$, and this cycle has no entry in $M$. Let $v\in M$ be a vertex on this cycle. Then $\textup{C}^{*}(G)_{\Omega(M)}\mathrel{\triangleleft}\textup{C}^{*}(G)$ is a prime ideal in $\textup{C}^{*}(G)$ and the subquotient
$\textup{C}^{*}(G)|_{M}\mathrel{\vcentcolon=}\textup{C}^{*}(G)_{\langle\Omega(M)\cup\{v\}\rangle}/\textup{C}^{*}(G)_{\Omega(M)}$
is gauge-simple and Morita–Rieffel equivalent to $\mathrm{C}(\mathbb{T})$. More precisely, the image of the corner $p_{v}\textup{C}^{*}(G)p_{v}$ in $\textup{C}^{*}(G)/\textup{C}^{*}(G)_{\Omega(M)}$ is isomorphic to $\mathrm{C}(\mathbb{T})$, generated by the image of $t_{e_{1}}t_{e_{2}}\dotsm t_{e_{k}}$, which is unitary. We call $\textup{C}^{*}(G)|_{M}$ the gauge-simple subquotient of $\textup{C}^{*}(G)$ associated to $M\in\mathcal{M}_{\tau}(G)$.
The corner $p_{v}(\textup{C}^{*}(G)/\textup{C}^{*}(G)_{\Omega(M)})p_{v}$ is Morita–Rieffel equivalent to the ideal in $\textup{C}^{*}(G)/\textup{C}^{*}(G)_{\Omega(M)}$ generated by the image of $p_{v}$. This ideal is gauge-invariant and corresponds to the smallest hereditary and saturated subset of $V$ containing $\Omega(M)$ and $v$. Then it must be $\textup{C}^{*}(G)|_{M}$.
Now we return to the situation of Theorem 8.7. That is, $G$ and $G^{\prime}$ are countable graphs, $A=\textup{C}^{*}(G)$, $B=\textup{C}^{*}(G^{\prime})$, and there are an order isomorphism $\psi\colon\mathbb{I}_{\mathrm{p}}(G)\cong\mathbb{I}_{\mathrm{p}}(A)\xrightarrow{\sim}\mathbb{I}_{\mathrm{p}}(B)\cong\mathbb{I}_{\mathrm{p}}(G^{\prime})$ and diagram isomorphisms
$\varphi_{0}^{+}\colon\textup{K}_{0}^{+}(A,\mathbb{I}_{\mathrm{p}})$ $\xrightarrow{\sim}\psi^{*}\textup{K}_{0}^{+}(B,\mathbb{I}_{\mathrm{p}}),$
$\varphi_{1}\colon\textup{K}_{1}(A,\mathbb{I}_{\mathrm{p}})$ $\xrightarrow{\sim}\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}}).$
These isomorphisms induce a monoid isomorphism $\varphi_{0}\colon\mathcal{P}(A)\xrightarrow{\sim}\mathcal{P}(B)$. We assume that the obstruction class of $(\varphi_{0},\varphi_{1})$ vanishes.
Lemma 9.4.
The order isomorphism $\psi\colon\mathbb{I}_{\mathrm{p}}(G)\xrightarrow{\sim}\mathbb{I}_{\mathrm{p}}(G^{\prime})$ extends uniquely to an order isomorphism $\bar{\psi}\colon\mathbb{I}(G)=\mathbb{I}_{\mathrm{p}}^{\infty}(G)\xrightarrow{\sim}\mathbb{I}_{\mathrm{p}}^{\infty}(G^{\prime})=\mathbb{I}(G^{\prime})$. The extension maps the subsets $\mathcal{M}_{\gamma}(G)$ and $\mathcal{M}_{\tau}(G)$ bijectively onto $\mathcal{M}_{\gamma}(G^{\prime})$ and $\mathcal{M}_{\tau}(G^{\prime})$, respectively.
Proof.
As in the proof of Proposition 7.3, the extension $\bar{\psi}$ exists by Lemma 7.2. The order isomorphism $\bar{\psi}$ restricts to a bijection between the prime ideals. Thus it induces a bijection $\psi\colon\mathcal{M}(G)\xrightarrow{\sim}\mathcal{M}(G^{\prime})$. We must show that $M\in\mathcal{M}(G)$ belongs to $\mathcal{M}_{\tau}(G)$ if and only if $\psi(M)\in\mathcal{M}(G^{\prime})$ belongs to $\mathcal{M}_{\tau}(G^{\prime})$. To do this, we characterise through the diagrams $\textup{K}_{0}^{+}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})$ and $\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}})$ when $M\in\mathcal{M}(G)$ belongs to $\mathcal{M}_{\tau}(G)$. First, Remark 4.17 shows that $\varphi_{0}^{+}$ and $\varphi_{1}$ extend to isomorphisms of diagrams $\textup{K}_{0}^{+}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}}^{\infty})\xrightarrow{\sim}\textup{K}_{0}^{+}(\textup{C}^{*}(G^{\prime}),\mathbb{I}_{\mathrm{p}}^{\infty})$ and $\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I}_{\mathrm{p}}^{\infty})\xrightarrow{\sim}\textup{K}_{1}(\textup{C}^{*}(G^{\prime}),\mathbb{I}_{\mathrm{p}}^{\infty})$. Theorem 5.3 identifies $\mathbb{I}_{\mathrm{p}}^{\infty}(\textup{C}^{*}(G))$ with $\mathbb{I}(G)$, the lattice of hereditary saturated subsets.
Let $M\in\mathcal{M}(G)$. Then $H\mathrel{\vcentcolon=}\Omega(M)=V\setminus M\in\mathbb{I}(G)$. If $M\in\mathcal{M}_{\tau}(G)$, then we define $H_{2}\mathrel{\vcentcolon=}\langle\Omega(M)\cup\{v\}\rangle\in\mathbb{I}(G)$ as in Proposition 9.3. Then $H\subsetneq H_{2}$. If $K\in\mathbb{I}(G)$ also satisfies $H\subsetneq K$, then $K\cap H_{2}=H$ or $K\cap H_{2}=H_{2}$ because $K\cap H_{2}\in\mathbb{I}(G)$ and the subquotient $\textup{C}^{*}(G)|_{M}\mathrel{\vcentcolon=}\textup{C}^{*}(G)_{H_{2}}/\textup{C}^{*}(G)_{H}$ has no gauge-invariant ideals. The case $K\cap H_{2}=H$ is impossible because $H$ is prime. It follows that $H_{2}\in\mathbb{I}(G)$ is the unique minimal element of the set $\{K\in\mathbb{I}(G)\,{:}\,\mathopen{}H\subsetneq K\}$. Therefore, the existence of such a unique minimal element $H_{2}$ is necessary for $M\in\mathcal{M}_{\tau}(G)$. This property is preserved by order isomorphisms. We assume from now on that such an $H_{2}$ exists and, in this generality, we define
$\textup{C}^{*}(G)|_{M}\mathrel{\vcentcolon=}\textup{C}^{*}(G)_{H_{2}}/\textup{C}^{*}(G)_{H}.$
A quotient of a graph $\textup{C}^{*}$-algebra by a gauge-invariant ideal is again a graph $\textup{C}^{*}$-algebra (see [Bates-Pask-Raeburn-Szymanski:Row_finite]*Thereom 4.1), and the action induced by the gauge action is the usual gauge action. Any gauge-invariant ideal in a graph $\textup{C}^{*}$-algebra is again isomorphic to a graph $\textup{C}^{*}$-algebra by a gauge-equivariant isomorphism (see [Vas:Graded_ideal_Leavitt]). Therefore, $\textup{C}^{*}(G)|_{M}$ is $\mathbb{T}$-equivariantly isomorphic to the graph $\textup{C}^{*}$-algebra of some graph $G_{M}$. This subquotient of $\textup{C}^{*}(G)$ has no gauge-invariant ideals because $H_{2}$ is the unique minimal element of $\{K\in\mathbb{I}(G)\,{:}\,\mathopen{}H\subsetneq K\}$. Equivalently, the graph $G_{M}$ is cofinal. Now there are three possible cases:
  1. (1)
    $G_{M}$ has no cycles; then $\textup{C}^{*}(G_{M})$ is a simple AF-algebra;
  2. (2)
    $G_{M}$ has more than one simple cycle; then $\textup{C}^{*}(G_{M})$ is simple purely infinite;
  3. (3)
    $G_{M}$ contains exactly one simple cycle.
In the third case, the unique simple cycle cannot have an entry because then $\textup{C}^{*}(G_{M})$ would be simple by [Bates-Hong-Raeburn-Szymanski:Ideal_structure]*Proposition 5.1, and then we must be in one of the first two cases, by the well known dichotomy for simple graph $\textup{C}^{*}$-algebras. So $M\in\mathcal{M}_{\tau}(G)$ if and only if $\textup{C}^{*}(G_{M})$ is Morita–Rieffel equivalent to $\mathrm{C}(\mathbb{T})$.
The graph $G_{M}$ is described explicitly in [Bates-Pask-Raeburn-Szymanski:Row_finite] and [Vas:Graded_ideal_Leavitt]. Its vertex set is obtained from $M=H_{2}\setminus H$ by adding certain vertices. An inspection shows that the projections $p_{v}$ for the added vertices are all Murray–von Neumann equivalent to projections from vertices in $M$. Therefore, Theorem 3.1 implies that any projection in $\textup{C}^{*}(G_{M})$ lifts to a projection in $\textup{C}^{*}(G)_{H_{2}}$. That is, the map $\textup{K}_{0}^{+}(\textup{C}^{*}(G)_{H_{2}})\to\textup{K}_{0}^{+}(\textup{C}^{*}(G_{M}))$ is surjective. This remains so on the Grothendieck groups. Therefore, the K-theory long exact sequence for the $\textup{C}^{*}$-algebra extension $\textup{C}^{*}(G)_{H}\rightarrowtail\textup{C}^{*}(G)_{H_{2}}\twoheadrightarrow\textup{C}^{*}(G)|_{M}$ looks as follows:
[…]
If $M\in\mathcal{M}_{\tau}(G)$, then $\textup{C}^{*}(G)|_{M}$ is Morita–Rieffel equivalent to $\mathrm{C}(\mathbb{T})$ by Proposition 9.3 and therefore $\textup{K}_{0}^{+}(\textup{C}^{*}(G)|_{M})\cong\mathbb{N}$ and $\textup{K}_{1}(\textup{C}^{*}(G)|_{M})\cong\mathbb{Z}$. In contrast, if $M\not\in\mathcal{M}_{\tau}(G)$, then either $\textup{C}^{*}(G)|_{M}$ is purely infinite, so that $\textup{K}_{0}^{+}(\textup{C}^{*}(G)|_{M})$ is a group, or $\textup{C}^{*}(G)|_{M}$ is AF, so that $\textup{K}_{1}(\textup{C}^{*}(G)|_{M})\cong 0$. Now we check that these three cases can be distinguished looking only at the maps
$\textup{K}_{0}^{+}(\textup{C}^{*}(G)_{H})\to\textup{K}_{0}^{+}(\textup{C}^{*}(G)_{H_{2}}),\qquad\textup{K}_{i}(\textup{C}^{*}(G)_{H})\to\textup{K}_{i}(\textup{C}^{*}(G)_{H_{2}})$
for $i=0,1$ induced by the inclusion $H\subsetneq H_{2}$. Indeed, if $M\in\mathcal{M}_{\tau}(G)$, then the quotient of $\textup{K}_{0}^{+}(\textup{C}^{*}(G)_{H_{2}})$ by the image of $\textup{K}_{0}^{+}(\textup{C}^{*}(G)_{H})$ cannot be a group because it maps to the monoid $\textup{K}_{0}^{+}(\textup{C}^{*}(G)|_{M})\cong\mathbb{N}$, which is not a group, and it cannot happen that the map $\textup{K}_{0}(\textup{C}^{*}(G)_{H})\to\textup{K}_{0}(\textup{C}^{*}(G)_{H_{2}})$ is injective and $\textup{K}_{1}(\textup{C}^{*}(G)_{H})\to\textup{K}_{1}(\textup{C}^{*}(G)_{H_{2}})$ is surjective, because that would force $\textup{K}_{1}(\textup{C}^{*}(G)|_{M})$ to vanish. If $M\not\in\mathcal{M}_{\tau}(G)$ and $\textup{C}^{*}(G)|_{M}$ is purely infinite simple, then the quotient of $\textup{K}_{0}^{+}(\textup{C}^{*}(G)_{H_{2}})$ by the image of $\textup{K}_{0}^{+}(\textup{C}^{*}(G)_{H})$ is a group. And if $M\not\in\mathcal{M}_{\tau}(G)$ and $\textup{C}^{*}(G)|_{M}$ is AF, then the map $\textup{K}_{0}(\textup{C}^{*}(G)_{H})\to\textup{K}_{0}(\textup{C}^{*}(G)_{H_{2}})$ is injective and $\textup{K}_{1}(\textup{C}^{*}(G)_{H})\to\textup{K}_{1}(\textup{C}^{*}(G)_{H_{2}})$ is surjective. Thus we can read off from $\mathbb{I}(G)$ and the diagrams $\textup{K}_{0}^{+}(\textup{C}^{*}(G),\mathbb{I})$ and $\textup{K}_{1}(\textup{C}^{*}(G),\mathbb{I})$ whether or not $M\not\in\mathcal{M}_{\tau}(G)$. Since these invariants do not distinguish our two graphs, it follows that the bijection $\psi\colon\mathcal{M}(G)\xrightarrow{\sim}\mathcal{M}(G^{\prime})$ maps $\mathcal{M}_{\tau}(G)$ onto $\mathcal{M}_{\tau}(G^{\prime})$. ∎
Theorem 9.6.
Let $G$ and $G^{\prime}$ be countable graphs. Define $A\mathrel{\vcentcolon=}\textup{C}^{*}(G)$, $B\mathrel{\vcentcolon=}\textup{C}^{*}(G^{\prime})$. Let
$\psi\colon\mathbb{I}_{\mathrm{p}}(G)\cong\mathbb{I}_{\mathrm{p}}(A)$ $\xrightarrow{\sim}\mathbb{I}_{\mathrm{p}}(B)\cong\mathbb{I}_{\mathrm{p}}(G^{\prime})$
be an order isomorphism and let
$\varphi_{0}^{+}\colon\textup{K}_{0}^{+}(A,\mathbb{I}_{\mathrm{p}})$ $\xrightarrow{\sim}\psi^{*}\textup{K}_{0}^{+}(B,\mathbb{I}_{\mathrm{p}}),$
$\varphi_{1}\colon\textup{K}_{1}(A,\mathbb{I}_{\mathrm{p}})$ $\xrightarrow{\sim}\psi^{*}\textup{K}_{1}(B,\mathbb{I}_{\mathrm{p}})$
be diagram isomorphisms. These isomorphisms induce a monoid isomorphism $\varphi_{0}\colon\mathcal{P}(A)\xrightarrow{\sim}\mathcal{P}(B)$. Assume that the obstruction class of $(\varphi_{0},\varphi_{1})$ vanishes. Then there is an $A$-$B$-correspondence $\mathcal{H}_{AB}$ that induces the isomorphisms $\varphi_{0}^{+}$ and $\varphi_{1}$ and that induces a Morita–Rieffel equivalence between the gauge-simple subquotients of $A$ and $B$ associated to $M$ and $\psi(M)$ for all $M\in\mathcal{M}_{\tau}(G)$. In addition, there is a $B$-$A$-correspondence $\mathcal{H}_{BA}$ that is inverse to $\mathcal{H}_{AB}$ up to homotopy, and such that all the correspondences in the homotopies between the composite correspondences $\mathcal{H}_{AB}\otimes_{B}\mathcal{H}_{BA}$ and $\mathcal{H}_{BA}\otimes_{A}\mathcal{H}_{AB}$ and the identity correspondences induce the identity correspondences on the gauge-simple subquotients associated to all $M\in\mathcal{M}_{\tau}(G)$, and similarly for $\mathcal{H}_{BA}\otimes_{A}\mathcal{H}_{AB}$.
Proof.
First, we fix $M\in\mathcal{M}_{\tau}(G)$ and write it as $M=H_{2}\setminus H$ with hereditary and saturated subsets $H$ and $H_{2}$ as above. Let $H_{2}^{\prime}\mathrel{\vcentcolon=}\psi(H_{2})$ and $H^{\prime}\mathrel{\vcentcolon=}\psi(H)$ in $\mathbb{I}(G^{\prime})$, and let $M^{\prime}\mathrel{\vcentcolon=}H_{2}^{\prime}\setminus H^{\prime}$. Then $M^{\prime}\in\mathcal{M}_{\tau}(G^{\prime})$ by Lemma 9.4.
Theorem 8.7 shows that there is a $\textup{C}^{*}(G)$-$\textup{C}^{*}(G^{\prime})$-correspondence $\mathcal{H}_{AB}$ that induces the maps $\psi$, $\varphi_{0}^{+}$ and $\varphi_{1}$ on our invariants. In particular, this correspondence induces correspondences between the gauge-invariant ideals $\textup{C}^{*}(G)_{H}$ and $\textup{C}^{*}(G^{\prime})_{\psi(H)}$ for all $H\in\mathbb{I}_{\mathrm{p}}(G)$; this extends automatically to $H\in\mathbb{I}(G)$. And then it also induces a proper correspondence $\mathcal{H}_{AB,M}$ between the subquotients $\textup{C}^{*}(G)|_{M}$ and $\textup{C}^{*}(G^{\prime})|_{M^{\prime}}$. Since our correspondence is a homotopy equivalence that preserves the gauge-invariant ideals, it induces an isomorphism on $\textup{K}_{*}(\textup{C}^{*}(G)_{H})$ and $\textup{K}_{*}(\textup{C}^{*}(G)_{H_{2}})$ in (9.5). By the Five Lemma, it induces an isomorphism $\textup{K}_{*}(\textup{C}^{*}(G)|_{M})\xrightarrow{\sim}\textup{K}_{*}(\textup{C}^{*}(G^{\prime})|_{M^{\prime}})$ as well. In addition, this isomorphism on $\textup{K}_{0}(\textup{C}^{*}(G)|_{M})\cong\mathbb{Z}$ maps the positive cone $\mathbb{N}$ into itself. Thus it is the identity map.
Let $v\in M$ and $v^{\prime}\in M^{\prime}$ be vertices on the unique cycles in $M$ and $M^{\prime}$, respectively. The image of $p_{v}$ generates $\textup{K}_{0}^{+}(\textup{C}^{*}(G)|_{M})$, and any other idempotent that generates $\textup{K}_{0}^{+}(\textup{C}^{*}(G)|_{M})$ is Murray–von Neumann equivalent to it because $\textup{C}^{*}(G)|_{M}$ is Morita equivalent to $\mathrm{C}(\mathbb{T})$ by Proposition 9.3 and $\mathcal{P}(\mathrm{C}(\mathbb{T}))\cong\mathbb{N}$. Similarly, any generator of $\textup{K}_{0}^{+}(\textup{C}^{*}(G^{\prime})|_{M^{\prime}})$ is Murray–von Neumann equivalent to $p_{v^{\prime}}$ in $\textup{C}^{*}(G^{\prime})_{M^{\prime}}$, and the corner in $\textup{C}^{*}(G^{\prime})_{M^{\prime}}$ that it cuts out is isomorphic to $\mathrm{C}(\mathbb{T})$ as well. Since the projective module $p_{v}\mathcal{H}_{AB,M}$ over $\textup{C}^{*}(G^{\prime})_{M^{\prime}}$ is a generator of $\mathcal{P}(\textup{C}^{*}(G^{\prime})|_{M^{\prime}})$, it is isomorphic to $\textup{C}^{*}(G^{\prime})|_{M^{\prime}}p_{v^{\prime}}$. We fix such an isomorphism $W\colon p_{v}\mathcal{H}_{AB,M}\xrightarrow{\sim}\textup{C}^{*}(G^{\prime})|_{M^{\prime}}p_{v^{\prime}}$. Let $z\in\mathrm{C}(\mathbb{T})\cong p_{v}\textup{C}^{*}(G)|_{M}p_{v}\cong\mathrm{C}(\mathbb{T})$ be the canonical unitary generator, and also write $z$ for the operator of left multiplication by $z$ on $p_{v}\mathcal{H}_{AB,M}$. Then $WzW^{*}$ is a unitary in $\mathbb{K}(\textup{C}^{*}(G^{\prime})|_{M^{\prime}}p_{v^{\prime}})\cong p_{v^{\prime}}\textup{C}^{*}(G^{\prime})|_{M^{\prime}}p_{v^{\prime}}\cong\mathrm{C}(\mathbb{T})$. Since $\mathcal{H}_{AB,M}$ induces an isomorphism on $\textup{K}_{1}$, the image of $WzW^{*}$ in $\textup{K}_{1}(\mathrm{C}(\mathbb{T}))\cong\mathbb{Z}$ is a generator. Since $\mathrm{C}(\mathbb{T})$ is $\textup{K}_{1}$-bijective, say, by Theorem 6.9, $WzW^{*}$ is homotopic to either $z$ or $z^{-1}$. Then we may write $WzW^{*}=\bar{\zeta}\cdot z^{\pm 1}$ with a unitary $\bar{\zeta}$ in $p_{v^{\prime}}\textup{C}^{*}(G^{\prime})|_{M^{\prime}}p_{v^{\prime}}$ that is homotopic to the unit. Then the unitary $W^{*}\bar{\zeta}^{*}W\in\mathbb{K}(p_{v}\mathcal{H}_{AB,M})$ is homotopic to $1$ as well. Therefore, we may lift $W^{*}\bar{\zeta}^{*}W$ to a unitary $\Upsilon_{v}\in\mathbb{K}(p_{v}\mathcal{H}_{AB})|_{H_{2}}$ homotopic to $1$ along the quotient map to $\mathbb{K}(p_{v}\mathcal{H}_{AB,M})$. When we multiply the unitary $U_{v}$ in our description of $\mathcal{H}_{AB}$ by $\Upsilon_{v}$, we do not change the homotopy class of $\mathcal{H}_{AB}$. Thus the induced maps on $\textup{K}_{0}^{+}$ and $\textup{K}_{1}$ also remain the same. By Proposition 9.3, the element $z\in\mathrm{C}(\mathbb{T})\subseteq p_{v}\textup{C}^{*}(G|_{M})p_{v}$ is the product of the edge generators in the cycle in $M$ based at $v$. So it contains a matrix coefficient of $U_{v}$ exactly once. Therefore, after multiplying $U_{v}$ by $\Upsilon_{v}$, we have arranged that the induced map
(9.7) $\mathrm{C}(\mathbb{T})\cong p_{v}\textup{C}^{*}(G)|_{M}p_{v}\to\mathbb{K}(\mathcal{H}_{AB,M})\cong p_{v^{\prime}}\textup{C}^{*}(G^{\prime})|_{\psi(M)}p_{v^{\prime}}\cong\mathrm{C}(\mathbb{T})$
is the isomorphism induced by $z\mapsto z^{\pm 1}$. Thus $\mathcal{H}_{AB,M}$ has become a Morita equivalence bimodule.
The only ingrediet of $\mathcal{H}_{BA}$ that we have changed to achieve this is the unitary $U_{v}$ for one vertex $v\in M$. These changes for different $M\in\mathcal{M}_{\tau}(G)$ do not interfere with each other. Therefore, we may arrange for $\mathcal{H}_{AB,M}$ to become a Morita equivalence bimodule for all $M\in\mathcal{M}_{\tau}(G)$ simultaneously.
Lemma 8.6 gives a $B$-$A$-correspondence $\mathcal{H}_{BA}$ that is inverse to $\mathcal{H}_{AB}$ up to homotopy. Even more, the homotopies that are implicit here automatically preserve all gauge-invariant ideals because equivalent projections generate the same ideal. Therefore, for each $M\in\mathcal{M}_{\tau}(G)$, there are induced correspondences $\mathcal{H}_{AB,M}$ and $\mathcal{H}_{BA,M}$ between $\textup{C}^{*}(G)|_{M}$ and $\textup{C}^{*}(G^{\prime})|_{\psi(M)}$ that are inverse to each other up to homotopy.
In the same way as above, we may replace $\mathcal{H}_{BA}$ by a homotopic proper correspondence that induces Morita equivalence bimodules on the subquotients for all $M^{\prime}\in\mathcal{M}_{\tau}(G^{\prime})$. Even more, we arranged both $\mathcal{H}_{AB}$ and $\mathcal{H}_{BA}$ to map the preferred generator $z$ of $\mathrm{C}(\mathbb{T})$ to $z$ or $z^{-1}$. Hence so do the composites $\mathcal{H}_{AB}\otimes_{B}\mathcal{H}_{BA}$ and $\mathcal{H}_{BA}\otimes_{A}\mathcal{H}_{AB}$. Since these composites are homotopic to the identity correspondence, the map $z\mapsto z^{-1}$ cannot occur here. Therefore, we know that the composite correspondences $\mathcal{H}_{AB}\otimes_{B}\mathcal{H}_{BA}$ and $\mathcal{H}_{BA}\otimes_{A}\mathcal{H}_{AB}$ both induce the identity bimodules on the subquotients for elements of $\mathcal{M}_{\tau}(G)$. For each $M\in\mathcal{M}_{\tau}(G)$, the homotopy between $\mathcal{H}_{AB}\otimes_{B}\mathcal{H}_{BA}$ and the identity correspondence on $A$ induces a homotopy of -homomorphisms $\mathrm{C}(\mathbb{T})\to\mathrm{C}(\mathbb{T})$ of the form $z\mapsto\bar{\zeta}_{t}\cdot z$ for some continuous path of unitaries $\bar{\zeta}_{t}$, where we already arranged that $\bar{\zeta}_{t}=1$ for $t=0,1$. We may lift these unitaries as above and correct the $\textup{C}^{*}(G)$-$\textup{C}^{*}(G)$-correspondences $(\mathcal{H}_{t})_{t\in[0,1]}$ in the homotopy, so that $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ remain the same and each $\mathcal{H}_{t}$ induces the identity map on the corner $\mathrm{C}(\mathbb{T})\cong p_{v}\textup{C}^{*}(G)|_{M}p_{v}$ for all $M\in\mathcal{M}_{\tau}(G)$ and some $v\in M$ in the unique cycle in $M$. This yields a new homotopy between $\mathcal{H}_{AB}\otimes_{B}\mathcal{H}_{BA}$ and the identity correspondence, with the extra property that each correspondence in the homotopy induces the identity Hilbert bimodule on $\textup{C}^{*}(G)|_{M}$ for all $M\in\mathcal{M}_{\tau}(G)$. The same change of homotopy is possible for the other composite $\mathcal{H}_{BA}\otimes_{A}\mathcal{H}_{AB}$. ∎